### v1

Abstract
This position paper examines the use of Large Language Models (LLMs) in social simulation, analyzing both their potential and their limitations from a computational social science perspective. The first part reviews recent findings on the ability of LLMs to replicate key aspects of human cognition, including Theory of Mind reasoning and social inference, while also highlighting significant limitations such as cognitive biases, lack of true understanding, and inconsistencies in behavior. The second part surveys emerging applications of LLMs in multi-agent simulation frameworks, focusing on system architectures, scale, and validation strategies. Notable projects such as Generative Agents (Smallville) and AgentSociety are discussed in terms of their design choices, empirical grounding, and methodological innovations. Particular attention is given to the challenges of behavioral fidelity, calibration, and reproducibility in large-scale LLM-driven simulations. The final section distinguishes between contexts where LLMs, like other “black-box” systems, offer direct value—such as interactive simulations and serious games—and those where their use is more problematic, notably in explanatory or predictive modeling. The paper concludes by advocating for hybrid approaches that integrate LLMs into traditional agent-based modeling platforms (GAMA, Netlogo, etc), enabling modelers to combine the expressive flexibility of language-based reasoning with the transparency and analytical rigor of classical rule-based systems.
keywords: Large Language Models, Social Simulation, Agent-Based Modeling, Human Behavior Modeling
1 Introduction
Large Language Models (LLMs) have emerged as a transformative class of deep learning architectures capable of generating human-like text across a wide range of domains. Built upon the transformer architecture, which employs parallelized attention mechanisms rather than sequential processing, these models are pre-trained on massive corpora comprising web data, encyclopedic content, and code repositories (Vaswani 2017; Devlin 2019; Brown et al. 2020). Through unsupervised or self-supervised learning, LLMs acquire syntactic, semantic, and pragmatic language competencies, enabling them to perform a broad array of downstream tasks such as question answering, summarization, translation, and code generation (Tonmoy 2024). Importantly, the magnitude and diversity of these training corpora mean that LLMs have effectively absorbed an immense statistical imprint of human language use across cultures, domains, and contexts. This goes far beyond learning grammar or syntax: LLMs encode traces of how people reason, argue, empathize, hesitate, and make decisions — all embedded in the linguistic fabric of their training data. Subsequently, these models are fine-tuned via instruction tuning, a supervised process in which models learn to follow natural language commands across a wide range of tasks. This is typically achieved using curated instruction datasets (Sanh et al. 2021; Wei et al. 2021), sometimes augmented with human feedback to better align model outputs with user intent (Ouyang et al. 2022).
A critical area of current investigation concerns the capacity of LLMs to simulate human behavior. One classical benchmark in this domain is the Turing Test, proposed by Alan Turing in 1950 as the imitation game, which evaluates whether a machine can produce conversational behavior indistinguishable from that of a human interlocutor (Jones & Bergen 2025). Recent empirical findings suggest that state-of-the-art LLMs, notably LLaMa-3.1 and GPT-4.5, have demonstrated convincing performance on a three-party version of this test. In these evaluations, LLaMa-3.1 was perceived as human in 56% of cases, while GPT-4.5 achieved a 73% success rate (Jones & Bergen 2025). This represents a notable advancement over previous models, which had only succeeded under less rigorous two-party test conditions. These results underscore the impressive ability of modern LLMs to adopt complex linguistic styles and behavioral nuances, particularly when guided by carefully crafted persona prompts — a form of cognitive prompt engineering that remains a highly precise and non-trivial technique requiring expert design and task sensitivity.
The apparent success of LLMs in the Turing Test has naturally prompted speculation about their suitability for simulating human agents within computational social systems (Park 2023; Anthis et al. 2025; Kirkeby-Hinrup & Stenseke 2025). However, the capacity of LLMs to produce believable human-like dialogue is not indicative of genuine understanding or consciousness. Rather, their performance reflects advanced statistical pattern recognition — the ability to mimic linguistic structures convincingly — without necessarily modeling the underlying cognitive or emotional states of a human agent (Jones & Bergen 2025). The tendency of human evaluators to anthropomorphize fluent language exacerbates this illusion, as observers may attribute intentionality or mental states to outputs that are, in essence, probabilistically generated (Jones & Bergen 2025), something well-known in psychology as the intentional stance (Dennett 1989). This distinction between functional imitation and authentic comprehension is particularly salient for social simulation. If LLMs merely replicate surface-level features of human behavior, their utility for generating novel insights into social processes — rather than reproducing known patterns — is fundamentally limited. Moreover, their use as proxies for human agents in scientific studies raises questions about external validity, representational fidelity, and cognitive realism. These concerns constitute a central theme of this paper and inform our critical assessment of LLMs as generative agents in artificial societies.
This position paper proceeds as follows. The first section examines the psychological plausibility of LLMs, focusing on their performance in tasks designed to evaluate theory of mind, social reasoning, and behavioral consistency. The second section surveys recent efforts to embed LLMs in agent-based social simulations, analyzing architectural strategies, scalability, and validation methodologies. The third section discusses both the potential contributions and limitations of LLM-based simulations, with particular emphasis on issues such as bias, interpretability, computational cost, and appropriate domains of application. Finally, the last section outlines future directions, including the integration of LLMs into classical modeling environments (e.g., GAMA, NetLogo) to support hybrid models that combine the explanatory clarity of rule-based approaches with the generative flexibility of LLMs.
2 LLMs in Representing Human Behavior
A critical question in assessing the suitability of LLMs for social simulation lies in evaluating how accurately they capture human psychological, emotional, and behavioral traits. Trained on massive corpora encompassing encyclopedic texts, social media exchanges, literature, forums, and dialogic interactions, LLMs have absorbed an unprecedentedly broad cross-section of human communicative behavior. As a result, they encode statistical traces of how people speak, reason, interact, and respond across a wide range of contexts — from mundane routines to emotionally charged situations. These models are thus not merely text generators, but synthesizers of large-scale sociolinguistic regularities, capable of emulating human-like behavior patterns precisely because such patterns are deeply embedded in their training data. Crucially, this capacity does not stem from intentional modeling or experiential understanding, but from the ability to compress and recompose behavioral signals at scale. When prompted appropriately, LLMs can reconstruct belief systems, emotional tones, and social scripts that mirror those present in the corpus — not through explicit design, but as emergent outcomes of large-scale pattern recognition. This gives rise to a powerful illusion: the model appears to “understand,” when in fact it is statistically interpolating between trillions of human-generated expressions.
Thus, while LLMs can convincingly mimic language use and pass superficial behavioral tests, the extent to which their outputs align with actual human cognition, reasoning patterns, emotional expression, and social bias remains contested.For instance, recent studies, such as Molchanova et al. (2025), raise important concerns about the depth and consistency of LLMs’ psychological realism by exploring their potential to simulate stable and diverse personality traits. This section provides a structured review of the evidence regarding these capabilities and limitations.
2.1 Theory of Mind and Cognitive Reasoning Tasks
Theory of Mind (ToM) refers to the human capacity to attribute beliefs, intentions, desires, and knowledge to others. It is considered a foundational element of social cognition and typically emerges in early childhood. One of the canonical tests to assess ToM in developmental psychology is the Sally-Anne test (Baron-Cohen et al. 1985). In this task, a child observes a scene in which Sally places a marble in her basket and leaves the room. Anne then moves the marble to a box. Upon Sally’s return, the child is asked where she will look for her marble. A correct response—indicating Sally will look in the basket—demonstrates an understanding that Sally holds a false belief, which differs from the child’s own knowledge.
Recent studies have assessed the capacity of LLMs to pass such false-belief tasks. Kosinski (2023) tested multiple LLMs, including GPT-3.5 and GPT-4, on 40 ToM problems designed to mirror psychological benchmarks. While early models performed near chance levels, GPT-4 achieved over 75% accuracy, comparable to that of a 6 or 7-year-old child. Furthermore, GPT-4 demonstrated competence in second and third order ToM reasoning (e.g., “I believe that she thinks that he knows…”), with occasional success even at fifth and sixth-order levels, surpassing typical adult human performance in structured tasks.
The robustness of these findings has been explored in follow-up work. Jaiswal (2024) tested LLMs on modified ToM tasks involving less prototypical narratives—such as a cat hiding a key or a user renaming folders offline—and found a significant drop in model performance. Performance was highly sensitive to phrasing; LLMs responded more accurately to questions using cognitive verbs (e.g., “think”) than to those phrased with action-oriented verbs (e.g., “look”), suggesting shallow linguistic pattern recognition rather than deep inference.
In addition to false-belief tasks, LLMs have been evaluated on their ability to interpret irony, indirect speech acts, and social narratives. Strachan et al. (2024) found that GPT-4 matched or exceeded average human performance in recognizing indirect requests, detecting sarcasm, and attributing communicative intent in narrative contexts. However, these results vary significantly with prompt design and context framing, raising questions about generalizability.
2.2 Emotion Representation and Affective Simulation
Emotions play a crucial role in human decision-making and social interaction. Unlike rule-based agents or earlier AI systems, LLMs can generate emotionally appropriate language in many contexts (Lee et al. 2024). For example, given a prompt describing a tragic event, GPT-4 can produce empathetic responses and plausible emotional narratives. Nonetheless, this capacity is fundamentally linguistic. LLMs do not experience emotions nor maintain internal emotional states. Their affective responses are generated through statistical associations learned from text corpora containing human emotional expression.
Studies have attempted to evaluate LLMs’ performance on emotional recognition tasks. Lai et al. (2025) report that LLMs can often classify emotions expressed in text (e.g., anger, sadness) and match tones to situational prompts. Yet these capabilities depend on typical lexical cues and falter when faced with indirect or culturally nuanced expressions. Emotional modeling in LLMs remains limited to surface-level mimicry, lacking integration with motivational systems or physiological grounding, both of which are central to human affective behavior.
2.3 Bias, Representational Gaps, and Cognitive Illusions
Because LLMs are trained on large-scale human-generated data, they inevitably inherit societal biases embedded in those texts. These include gender, racial, and cultural stereotypes. Studies have shown that LLMs reproduce associations consistent with biased representations: for instance, associating certain professions with male pronouns or producing racially skewed descriptions of criminal behavior(Santurkar 2025). In addition, structural imbalances in data availability mean that voices from underrepresented linguistic or social groups are poorly modeled, resulting in inequitable performance across different populations (Kotek et al. 2024).
Beyond representational biases, LLMs also exhibit patterns consistent with cognitive biases observed in humans.(Kim et al. 2025) found that LLMs replicate anchoring effects, status quo bias, framing effects, and recency bias when posed with structured reasoning tasks. Alarmingly, the tendency to apply ”chain-of-thought” prompting—a method designed to enhance logical reasoning—did not consistently reduce these biases. In some instances, reasoning chains amplified them, indicating that LLM ”reasoning” may reinforce pre-trained associations rather than correct them.
Efforts to debias LLMs using prompt engineering or few-shot learning strategies have shown promise but remain partial solutions. Schramowski (2024) argues that mitigating bias in LLMs requires not only refined prompting but also upstream control over training data composition and filtering mechanisms.
2.4 Limitations in Consistency, Hallucination, and Generalization
A pervasive challenge in LLM behavior is the issue of hallucination, where the model generates outputs that are linguistically fluent but factually incorrect or inconsistent with the input (Tonmoy 2024). This phenomenon undermines the credibility of LLM-generated reasoning and introduces substantial risks in simulation contexts. Combined with susceptibility to input phrasing and lack of state consistency over time, hallucination further limits the reliability of LLMs for modeling human behavior dynamically.
Another recurrent issue is inconsistency. Identical or nearly identical inputs can elicit divergent responses, even within the same session. This instability hinders reproducibility in simulated environments and challenges the interpretability of agent behavior. Wu et al. (2025) emphasizes that LLM agents, when evaluated across multiple instantiations, display lower variance and reduced behavioral diversity compared to human participants, while also amplifying social stereotypes.
2.5 Summary and Implications
The current landscape presents a paradox. LLMs exhibit impressive capabilities in tasks traditionally considered strong indicators of social cognition. They can perform multi-level mental state inference, recognize irony, and produce emotionally plausible responses. However, these abilities are brittle, highly sensitive to phrasing, and grounded in probabilistic mimicry rather than experiential or embodied cognition. Their outputs reflect statistical associations from training corpora, not inferential reasoning rooted in a coherent world model.
In light of these findings, LLMs offer valuable opportunities for simulating linguistic and behavioral patterns within constrained contexts. Yet, their application in scientific investigations of human psychology or complex social behavior should be approached with caution. Their limitations—opacity, hallucination, inconsistency, bias, and shallow understanding—necessitate robust validation, controlled environments, and clear demarcation between imitation and cognition. In the following section, we turn to how these capabilities are being harnessed in practical social simulation frameworks, and how model architecture and evaluation are being addressed in emerging systems.
3 LLMs in Social Simulation
The use of LLMs in social simulations builds upon years of research that have seen the application of machine learning techniques in multi-agent systems and, by extension, in agent-based models (Drogoul & Zucker 1998). It is therefore essential to understand whether LLMs represent a major breakthrough or, more modestly, a logical evolution of this partnership, and to determine if they share the same methodological challenges as previous approaches or introduce new ones. Building on the previous discussion regarding the representational capacities of LLMs in psychological and behavioral contexts, this section examines their application in the field of social simulation. We focus on how they are currently used to instantiate artificial societies, the architectural strategies adopted for agent based modeling, and the methodological challenges related to validation and scaling.
3.1 Foundational and Emerging Platforms
A foundational illustration of LLM-based social simulation is the Generative Agents framework introduced by Park (2023), widely known as the “Smallville” project. In this system, twenty-five agents inhabit a sandbox environment reminiscent of The Sims, where they autonomously engage in daily routines, form interpersonal relationships, and generate emergent collective behaviors. Each agent maintains a natural language memory stream, allowing it to reflect on past events, summarize experiences, and formulate plans. Remarkably, the simple assignment of a goal—such as organizing a party—was sufficient to initiate a chain of spontaneous, coordinated interactions among agents. Evaluation efforts focused primarily on assessing the believability and internal coherence of these emergent dynamics, rather than on their correspondence to empirical data.
More recent efforts have extended this line of research to address questions of scale, complexity, and empirical grounding. One prominent example is AgentSociety, a large-scale platform designed to simulate the dynamics of human societies with over 10,000 LLM-based agents (Piao et al. 2025). Each agent is instantiated using GPT-4 and equipped with cognitive modules for memory, goal management, decision-making, and social relationship tracking. These agents interact within rich environments encompassing spatial mobility, economic activity, social networks, and exogenous events such as natural disasters or political crises. AgentSociety has been applied to simulate phenomena including political polarization, rumor diffusion, and the societal impacts of universal basic income. Its validation strategy emphasizes empirical fidelity by replicating outcomes observed in behavioral experiments and large-scale social surveys.
Complementing these large-scale efforts, Simulate Anything offers a flexible platform for generating demographically diverse agent populations using LLMs, leveraging real-world user data for initialization and sampling (Wang 2025). The system supports scenario-specific tuning and hypothesis testing across domains such as politics, journalism, and economics. Its principal contribution lies in enabling heterogeneous role configurations and behavioral variability, though its validation methodologies are comparatively underdeveloped.
Another significant contribution in this domain is S3 (Social-network Simulation System), which specifically focuses on social network dynamics through LLM-empowered agents (Gao et al. 2025). S3 demonstrates particular strengths in simulating three pivotal aspects of human social behavior: emotion, attitude, and interaction behaviors. It leverages advanced prompt engineering and prompt tuning techniques to ensure authentic human-like agent behavior within social networks.
Several more recent platforms have further advanced the field in terms of architectural modularity and methodological rigor. GenSim, for instance, is a general-purpose simulation engine that supports tens of thousands of LLM agents (Tang et al. 2025). It introduces structured abstractions for social routines, long-term memory management, and adaptive behavior correction, all orchestrated by a dynamic scheduling system. Validation combines empirical replication, face validity assessments, and anomaly detection to ensure both plausibility and statistical consistency.
Another emerging contribution is AgentTorch, a modular and scalable framework that unifies large-scale agent-based modeling with differentiable programming and LLM-powered behavior synthesis (Chopra et al. 2023). Instead of relying on expensive real-time prompting for each agent, AgentTorch uses a small number of archetypal LLM agents to generate compact, reusable policy representations. These are then deployed across massive populations of lightweight agents in GPU-accelerated simulations. Whereas narrative-first systems emphasize fine-grained agent individuality, AgentTorch focuses on experimental control, statistical robustness, and scientific reproducibility, making it particularly suited for large-scale hypothesis testing and parameter exploration in computational social science.
SALLMA (Scalable Architecture for LLM Multi-Agent Applications) adopts a layered architecture, separating operational processes (intent formation, task execution, communication) from knowledge-level components (agent profiles, shared memory, workflows) (Becattini et al. 2025). This modular design facilitates agent specialization and integration with external tools via APIs. Its validation pipeline combines automated benchmarking with expert-in-the-loop review, aided by interactive dashboards and simulation playback for iterative refinement.
SocioVerse represents a shift toward population-scale calibration (Zhang et al. 2025b). It initializes agents from millions of real-world user profiles to mirror authentic distributions of attitudes, preferences, and social structures. The platform supports fine-grained tuning for scenario exploration, such as testing the effects of policy changes across diverse subpopulations. Its validation strategy emphasizes both empirical benchmarking and participatory evaluation, particularly focusing on the accurate representation of minority groups and on interactive adjustment of simulation parameters.
Similarly, the LLM-AIDSim framework integrates LLMs (Llama3:8b) into ABM for influence diffusion, enabling language-level responses and deeper insights into user agent interactions (Zhang et al. 2025a).
Collectively, these platforms reflect the rapid evolution of LLM-based social simulation toward greater scale, demographic realism, and methodological transparency. They also signal a growing reliance on hybrid validation strategies that combine quantitative benchmarking, expert evaluation, and dynamic, interactive diagnostics. Together, they mark a shift toward simulation systems that not only generate plausible behavior but also aspire to meet the standards of interpretability and scientific robustness necessary for credible computational social science.
3.2 Architectures for LLM-Based Multi-Agent Systems
3.2.1 Agent Representation and Cognitive Modules
LLM-based multi-agent models generally instantiate each agent as a distinct LLM instance or session using prompt engineering. Architectures typically involve three interconnected modules: a memory system that encodes and retrieves past experiences; a reflection and summarization layer that condenses observations into mental models; and a planning component that uses these reflections to generate actions (Park 2023; Li et al. 2024). These modules are often implemented as chained or nested prompts, with each step leveraging the LLM’s generative capacity to simulate cognitive processes such as recall, reasoning, and intention formation. For instance, the Humanized Agent-Based Models (h-ABM) framework, proposed by Atchade et al. (2024), outlines a modular architecture for integrating LLMs into ABMs to achieve more cognitively and emotionally realistic agents.
Memory Systems and Psychological Foundations
Many LLM-based agent architectures draw inspiration from cognitive psychology, particularly theories of episodic and semantic memory. For example, the memory systems in Generative Agents and AgentSociety are designed to mimic human-like recall, using vector databases or embedding-based retrieval to surface relevant past experiences. Some frameworks explicitly reference psychological models, such as the Atkinson-Shiffrin multi-store model or the concept of “theory of mind,” to inform the design of agent cognition and social reasoning (Li et al. 2024; Doran & et al. 2023). Reflection modules often prompt the LLM to summarize recent events, extract intentions, or infer the mental states of other agents, thus operationalizing constructs from social and developmental psychology.
Established cognitive architectures such as SOAR (Laird 2012) and ACT-R (Anderson et al. 2004) have provided theoretical foundations for agent design. SOAR, developed by John Laird, Allen Newell, and Paul Rosenbloom at Carnegie Mellon University (Laird 2012), embodies the problem space hypothesis, suggesting that all goal-oriented behavior can be characterized as a search in a problem space defined by states, operators, and goals. This cognitive architecture divides knowledge into procedural memory (skills and know-how), semantic memory (facts about the world), and episodic memory (memories of experiences). Similarly, ACT-R (Adaptive Control of Thought-Rational) (Anderson et al. 2004) posits that human cognition can be modeled through interconnected cognitive and perceptual modules that operate according to rational principles. These architectures inform the contemporary design of LLM agents by providing structured frameworks for organizing memory, decision-making, and learning.
Planning and Decision-Making
Planning modules in LLM-based multi-agent systems typically exploit the model’s capacity to generate contextually appropriate actions by conditioning outputs on agent-specific goals, environmental affordances, and social context. While a subset of frameworks integrates reinforcement learning or utility-based decision mechanisms to refine agent strategies, the majority rely on prompt engineering and scenario-specific instructions to elicit desired planning behavior. This approach emphasizes the structured design of inputs to steer agents’ reasoning, allowing for complex decision-making without the need for explicit reward modeling.
A hybrid approach is exemplified by the LGC-MARL framework proposed by Jia et al. (2025), which combines an LLM-based planner with graph-structured multi-agent reinforcement learning. In this architecture, the LLM is tasked with decomposing high-level instructions into executable subtasks, evaluating their coherence and feasibility, and formulating collaborative strategies among agents before final action selection. In contrast, more conventional social simulation systems rely predominantly on prompt-engineering strategies to define agent roles, goals, and reasoning pathways, typically without integrating reinforcement signals or dynamic adaptation mechanisms (Junprung 2023). These approaches underscore the current design trade-off between architectural complexity and ease of interpretability in the planning components of LLM-driven simulations.
Orchestration and Communication
In LLM-based multi-agent systems, a centralized orchestration layer is commonly adopted to manage simulation time-stepping and regulate agent interaction flows. Such architectures typically support inter-agent communication via various message-passing mechanisms, including direct messaging, simulated surveys, and public broadcasts.
Approaches to orchestration and communication vary considerably depending on the complexity and goals of the simulation. For example, platforms like Generative Agents and AgentSociety utilize orchestration layers to manage complex interactions among thousands of agents, often based on sequential flows or perception-reflection-action loops. Others, such as GenSim, integrate structured abstractions for social routines and a dynamic scheduling system. More advanced frameworks, like AgentTorch, move away from costly sequential dependencies by generating reusable policy representations from a small number of archetypal LLM agents, which are then deployed at scale on lightweight agents for GPU-accelerated simulations. SALLMA proposes a layered architecture, separating operational processes (intention formation, task execution) from knowledge-level components (agent profiles, shared memory, workflows), thereby facilitating agent specialization and integration with external tools via APIs. These platforms often support modular ”blocks,” which are reusable behavioral units that allow for systematic experimentation and increased extensibility. Integration with external APIs (e.g., web search, databases, calculators) is also crucial, as it enables agents to access additional information and tools, thereby augmenting their reasoning and decision-making capabilities (Chen et al. 2024b).
3.2.2 Specialization and Behavioral Diversity
LLM-based multi-agent system (MAS) frameworks are progressively integrating sophisticated specialization protocols, allowing individual agents to adopt roles characterized by unique behavioral repertoires. This specialization is typically realized through several mechanisms: the application of tailored prompt templates that guide an agent’s responses (e.g., (Wei et al. 2022)), which can enforce specific communication styles or decision-making heuristics; the implementation of role-specific memory structures that store relevant knowledge and past interactions (e.g., (Park 2023)), enabling agents to leverage historical data pertinent to their designated function; and scenario-driven fine-tuning processes that adapt an agent’s core model to specific operational contexts (e.g., (Ouyang et al. 2022)), thereby optimizing their performance for particular tasks or environments. Furthermore, some cutting-edge systems are exploring methods to inject personality traits or demographic attributes into agents. This can be achieved either by conditioning prompts with specific personality descriptors (e.g., ”act as a skeptical researcher”) or by sampling attributes from real-world survey distributions (e.g., age, occupation, cultural background), thereby enhancing the realism and diversity of agent behaviors and fostering more nuanced social dynamics (Wang et al. 2024; Chen et al. 2024b). These innovative techniques clearly enhance existing population synthesis algorithms (e.g., (Chapuis et al. 2022)) and enable researchers to explore complex emergent behaviors within agent populations, thereby gaining deeper insights into human-like interactions.
3.3 Validation Strategies and Data Sources
3.3.1 Empirical Validation and Benchmarking
The validation of LLM-based social simulations remains an active and evolving area of research. Among the most rigorous efforts, frameworks like AgentSociety adopt multi-faceted validation strategies that compare simulated outputs with real-world data. Empirical benchmarking typically involves aligning simulation results with social indicators such as voter turnout, economic behavior, or survey-derived attitudinal distributions from established instruments like the General Social Survey or the World Values Survey (Piao et al. 2025). In addition, some studies replicate classical experimental paradigms—such as public goods games or rumor propagation scenarios—within the simulation environment, and evaluate whether the emergent outcomes converge toward established empirical findings.
Longitudinal consistency is also a critical focus, particularly in simulations involving extended time horizons. Researchers analyze whether agent behavior remains stable and coherent across simulation runs, and whether emergent patterns are reproducible under minor changes to initial conditions or prompt phrasing (Wu et al. 2025). These approaches aim to assess not only face plausibility but also the robustness and generalizability of the simulated social dynamics.
3.3.2 Human-in-the-Loop Evaluation
Despite advances in automation, human judgment continues to play a crucial role in validating LLM-driven simulations. Expert evaluation by domain specialists is frequently used to assess the believability, internal coherence, and sociological plausibility of agent behaviors. Structured rating protocols are often employed to ensure inter-rater reliability, although such methods remain sensitive to subjective bias. Crowdsourced evaluation has also been adopted in some frameworks, wherein large groups of non-expert raters assess the realism or diversity of agent interactions.
Calibration against real-world survey data constitutes another widely used strategy. Here, demographic or attitudinal profiles from empirical sources are used to condition LLM agents, and human evaluators subsequently assess the extent to which the simulated agents reflect the intended population characteristics (Chen et al. 2024b). Nevertheless, recent findings suggest that human raters may be influenced by non-epistemic factors—such as assertiveness, linguistic fluency, or the perceived authority of LLM outputs—leading to potential overestimation of their factual or analytical accuracy (Chen et al. 2024a).
3.3.3 Specialized Validation Methods
In response to the methodological challenges of evaluating LLM-generated behavior, several specialized validation approaches have emerged. Face validation techniques rely on the visualization of agent behavior, often through animation or dashboards, to elicit expert feedback on whether the behavior appears theoretically plausible (Park 2023). Exploratory model behavior analysis is another technique, in which researchers systematically vary input conditions to test whether emergent behaviors align with theoretical expectations. Statistical correspondence testing has been used to quantitatively compare simulation outcomes to known empirical distributions or outputs from analytical models. Sensitivity analysis further complements these methods by evaluating the robustness of simulation results to perturbations in initial conditions and parameter values.
Sensitivity analysis further complements these methods by evaluating the robustness of simulation results to perturbations in initial conditions and parameter values. One concrete instantiation of this multi-level validation approach is the S3 system (Gao et al. 2025), which conducts evaluations at both individual and population levels using real-world social network data. Their validation strategy demonstrates promising accuracy in replicating observed social phenomena, including the propagation of information, attitudes, and emotions across network structures.
3.3.4 Validation Challenges and Open Issues
Despite these developments, substantial obstacles remain. One pervasive challenge is the convergence of LLMs toward an ”average persona” (Wu et al. 2025), which can suppress behavioral heterogeneity and thereby reduce the capacity of simulations to model complex social phenomena that rely on diversity of perspective and action. In many domains, the lack of clear empirical ground truth complicates efforts to validate simulated phenomena with sufficient rigor. Another critical issue concerns temporal instability—agents may exhibit behavioral drift over the course of long simulation runs, compromising reproducibility and longitudinal coherence (Wu et al. 2025).
Human evaluation, while informative, introduces its own set of biases. Recent findings suggest that human raters may be influenced by non-epistemic factors—such as assertiveness, linguistic fluency, or the perceived authority of LLM outputs—leading to potential overestimation of their factual or analytical accuracy (Steyvers et al. 2024).
Moreover, recent studies underscore the limitations of omniscient simulation settings, where agents are given access to complete information. While such configurations enhance performance, they are far removed from real-world conditions, which are typically characterized by asymmetrical or incomplete information (Zhou et al. 2024).
In light of these challenges, Wu et al. (2025) proposes a pragmatic framework for determining the conditions under which LLM-based social simulations can meaningfully advance scientific understanding. They argue that these simulations are most defensible when used to explore collective patterns rather than individual decision trajectories, when agent outputs align with population-level averages despite limited within-population variance, and when validation methods are carefully matched to the research objectives and simulation scale.
3.3.5 Data Sources and Comparative Benchmarks
The development of standardized benchmarks has been instrumental in advancing the empirical evaluation of LLM agents in social simulations. For instance, the SocialIQA dataset (Sap et al. 2019) evaluates an LLM’s capacity for common sense social reasoning by presenting situations involving physical, social, and emotional causality and testing whether the model can infer likely consequences. Another example is CrowS-Pairs (Nangia et al. 2020), which assesses social bias by contrasting sentence pairs that differ only by protected attributes (e.g., gender or race), allowing systematic measurement of discriminatory tendencies. Additionally, benchmarks like BiosBias (De Arteaga et al. 2019) evaluate demographic fairness by examining whether the model’s predictions or representations vary unjustly depending on demographic factors such as gender or ethnicity. While these datasets do not address agent consistency across sessions directly, related work such as Zhang et al. (2022) examines the consistency of language models by testing response variability to semantically equivalent paraphrased inputs, providing useful insights into reproducibility concerns in agent behavior.
Comparative studies between human and LLM behavior further inform the validation landscape. The EPITOME battery (Karra et al. 2024) includes a series of experimental tasks designed to test Theory of Mind capabilities in both humans and LLMs. Results suggest that while LLMs can often match or even exceed human-level performance on certain tasks, they also exhibit systematic limitations, particularly in pragmatic reasoning and inference under uncertainty.
Together, these empirical, human-centered, and dataset-driven approaches constitute an emerging ecosystem of validation methodologies. Nevertheless, their integration into a cohesive and principled framework remains an open area of research—one that is essential for ensuring the scientific credibility and social relevance of LLM-based simulations.
3.4 Implications and Methodological Cautions
Integrating LLMs into social simulation offers compelling opportunities: simulations are faster to build, more flexible in scope, and capable of generating richly expressive agent behavior. However, they also introduce epistemic risks. LLMs, trained to imitate plausible language rather than model causal processes, may yield emergent behaviors that are linguistically coherent but sociologically superficial or misleading. The gap between internal statistical inference and actual human cognition (as discussed in the previous section) persists in these settings and is amplified at scale. As a result, LLM-based social simulations may be most suitable for exploratory modeling, hypothesis generation, or interactive training environments—rather than confirmatory empirical research. Their use demands rigorous validation frameworks, careful scenario scoping, and, where appropriate, hybrid integration with traditional rule-based or equation-based models. Future work must focus on developing principled approaches for behavioral diversity injection, cross-run reproducibility, and modular interpretability to ensure these systems can meaningfully contribute to computational social science.
4 Discussion
4.1 Strengths of LLMs in Social Simulation
LLMs offer a compelling alternative to traditional human-participant studies, providing a scalable and cost effective means to explore social scenarios (Anthis et al. 2025; Piao et al. 2025). They enable the rapid simulation of thousands of interactions, facilitating the study of large-scale social dynamics that would be logistically impractical or prohibitively expensive with human subjects. This capability significantly reduces the resource burden associated with recruiting, compensating, and managing human participants, making complex social experiments more accessible (Anthis et al. 2025; Piao et al. 2025). Furthermore, unlike rigid rule-based simulations, LLMs can sometimes exhibit unexpected emergent behaviors that were not explicitly programmed, potentially offering novel perspectives into complex social phenomena (Anthis et al. 2025). This capacity for spontaneous, human-like interaction can lead to fresh insights into social dynamics that might be difficult to anticipate with more “theoretical” or constrained models.
While LLMs are fundamentally data-driven—learning behavioral patterns from large-scale corpora—whereas classical cognitive architectures are typically theory-driven, grounded in formal models derived from case studies and psychological theory, both approaches can be leveraged to support exploratory theory building in social simulation, i.e. understanding the range of possible effects under specific behavioral assumptions before costly real-world experiments are conducted. This makes them valuable for pilot studies where the objective is to uncover possibilities rather than to confirm hypotheses (Anthis et al. 2025). Researchers can then leverage LLMs to identify numerous ways in which social phenomena might manifest, exploring the space of potential outcomes without the need to avoid false positives, which is crucial in early-stage research (Wu et al. 2025). LLMs also enable researchers to conduct ethical investigations into sensitive real-world issues, such as social bias or psychological stress, without subjecting human participants to potentially harmful experimental conditions (Anthis et al. 2025). This provides a safe and controlled environment for exploring complex and ethically challenging scenarios.
The inherent natural language processing capabilities of LLMs allow generative agents to offer intuitive interfaces for human interaction, making simulations more flexible and their internal processing potentially more explainable (Li et al. 2024). This natural language interface simplifies the design and interpretation of agent behaviors, allowing researchers to define complex roles and interactions using human-readable prompts. For instance, LLMs can process and summarize large institutional documents, enabling their integration as proxies for organizational actors within socio-environmental simulations (Zeng et al. 2025).
Beyond these general benefits, LLMs offer tangible enhancements to established simulation frameworks. Consider, for example, a long-standing epidemiological simulation model like COMOKIT (Gaudou et al. 2020; Taillandier et al. 2024), designed to model disease spread and policy interventions. Traditionally, agent behaviors in COMOKIT are rule-based, limiting their adaptability to nuanced human responses or contextual variations. By integrating LLM-driven agents into such a framework, researchers could enable more realistic behavioral responses. For instance, instead of static compliance rates, LLM agents could interpret public health messages, react to evolving social norms, or even engage in simulated disinformation spread, reflecting real-world complexities. This allows for the exploration of rich, context-sensitive human dynamics – such as hesitancy towards vaccination campaigns or adherence to social distancing measures – which are challenging to capture with purely rule-based systems. LLMs’ ability to process and summarize large institutional documents (e.g., policy papers, news articles) further facilitates the definition of complex agent behaviors, enabling a deeper exploration of socio-environmental interactions and policy implications within existing, well-calibrated models.
Their ability to emulate context-sensitive reasoning and reflect demographically correlated response patterns highlights their potential for modeling institutional agency and culturally embedded decision-making processes. Such models can facilitate new forms of human–AI collaboration in complex domains requiring interpretability, adaptive reasoning, and integration across multiple knowledge sources (Zeng et al. 2025).
4.2 Limitations and Points of Vigilance
The ”black-box” nature of Large Language Models (LLMs), stemming from their intricate architectures with billions of parameters and complex non-linear transformations, creates significant challenges for understanding their internal workings and decision-making processes (Li 2024). In general, this lack of transparency poses substantial challenges for interpretability (understanding why an LLM arrives at a particular output becoming a speculative exercise), accountability, and trustworthiness, but it can also contribute to ”automation bias”. This bias manifests when modellers, particularly those lacking deep expertise in a given subject matter, excessively rely on LLM-generated information, often without critical evaluation. The perceived sophistication and apparent fluency of LLM outputs can effectively lead modelers to over-attribute accuracy and authority to the model, even when its underlying reasoning is obscure or flawed (Cummings 2012). So when LLMs are used to drive the behavior of agents in simulations, this ”black-box” issue makes it exceedingly difficult to verify the reliability and validity of the simulation’s results. For instance, if an agent’s decision is influenced by an LLM, and that decision is uninterpretable, assessing whether the simulated outcome accurately reflects reality or a desired theoretical model becomes nearly impossible. This can undermine the scientific rigor and credibility of the simulation. The unverified reliability of LLM-driven simulations also raises serious ethical and regulatory concerns, especially when the outputs of these simulations are intended to inform or directly influence decision-making processes (Li 2024). For instance, if a simulation guides policy decisions, resource allocation, or even emergency responses, the inability to fully understand and trust the LLM’s contribution to the simulation’s outcome introduces significant risks. Regulators and policymakers require assurances that the tools informing critical decisions are robust, transparent, and accountable, which the ”black-box” nature of LLMs directly contravenes. The potential for biased or inexplicable outputs from the LLM to cascade through the simulation, leading to unforeseen and potentially harmful real-world consequences, may necessitate a critical re-evaluation of how these agent-based models are integrated into decision-support systems.
In terms of ethical issues, a significant consequence of their “black-box” nature is that LLMs inherit and can perpetuate societal biases present in their vast training datasets (Santurkar 2025; Kim et al. 2025). These biases encompass explicit and implicit stereotypes related to race, gender, religion, and health, potentially leading to discriminatory outcomes and inconsistent user experiences (Santurkar 2025; Gallegos et al. 2024). Studies have found pervasive stereotype biases mirroring those in society across various social categories and stereotypes, demonstrating sizable effects on discriminatory decisions (Santurkar 2025). Furthermore, LLMs exhibit patterns resembling human cognitive biases (e.g., anchoring, framing, recency bias), which are often not directly visible and can implicitly mislead users (Schramowski 2024; Kim et al. 2025). Cognitive bias, defined as a systematic pattern of deviation from rationality in judgment, arises from information processing mechanisms and can lead to inconsistent decision-making (Schramowski 2024). Crucially, recent studies highlight that LLMs’ purported reasoning capabilities do not consistently prevent these cognitive biases; in some cases, applying reasoning even increases vulnerability to certain bias types (Kim et al. 2025). Bias persists in LLMs despite instruction tuning, and differences between model versions can lead to inconsistent user experiences and hidden harms in downstream applications (Gallegos et al. 2024).
Another bias that raises significant ethical issues is the convergence towards an “average persona”. LLMs fundamentally operate on statistical probabilities, predicting the ”next likely word” based on patterns in their training data (Wang et al. 2025). Their similarities to human language are purely functional, not indicative of deep cognitive understanding or conscious intent (Poeppel & Marantz 2025). While humans make decisions influenced by a rich tapestry of lived experiences, financial needs, and future goals, an LLM merely calculates the most expected response based on its training data (Wang et al. 2025) and LLMs therefore often converge towards an ”average persona”. This results in reduced behavioral variance and can inadvertently erase the characteristics of minority subgroups, leading to a lack of diversity in simulated populations (Wu et al. 2025; Gallegos et al. 2024). Empirical evidence shows LLMs producing highly uniform responses compared to the wider diversity observed in human choices, for instance, in economic games where LLMs consistently choose extreme values while humans exhibit a much wider range (Wu et al. 2025). This homogeneity significantly limits their capacity to accurately simulate complex social dynamics that inherently rely on individual and group heterogeneity (Wu et al. 2025). Furthermore, LLMs tend to produce narrower distributions of political opinions and may overrepresent opinions of wealthy, young, and politically liberal individuals in certain countries (Santurkar et al. 2023; Gover 2023). The way prompts are formulated can either reinforce or mitigate this convergence: generic prompts tend to elicit normative, majority-aligned responses, while more specific or context-rich prompts may allow for greater variation (Argyle et al. 2023). However, this variability remains contingent on the user’s intervention, not on an inherent capacity of the model to generate autonomous diversity. As a result, LLMs often fail to replicate the spontaneous variability and unpredictability that characterize real human societies—limiting their effectiveness for simulating pluralistic or emergent behaviors across diverse agent populations.
LLMs are prone to ”hallucination,” generating content that is baseless, untruthful, or diverges from factual reality, even while maintaining a coherent linguistic structure (Tonmoy 2024). Hallucinations can stem from flawed data sources or architectural biases (Li 2023; Huang et al. 2025; Gautam 2025). Crucially, when LLMs hallucinate, they often maintain an aura of authority and credibility, generating responses that appear coherent and well formed, which can mislead users, especially those lacking expertise (Simhi et al. 2025; Su 2024). Such deceptive coherence poses a significant risk in scientific applications where factual accuracy is paramount. Additionally, LLMs can exhibit ’inconsistency,’ producing varying or contradictory outputs for very similar (or even identical) inputs, which severely undermines their reliability and trustworthiness for scientific applications (Elazar et al. 2021). This inconsistency can leave lasting impressions of unreliability (Gao et al. 2022). LLMs demonstrate superior performance in unrealistic, ”omniscient” simulation settings where all agents have full information (Aher et al. 2023). However, they struggle significantly in scenarios that more accurately reflect real-world conditions characterized by information asymmetry, where agents possess incomplete or private knowledge (Zhou 2024). This fundamental mismatch with real-world information dynamics limits their applicability for accurate social prediction.
While social simulations, including the ones equipped with LLMs, offer cost-efficiency compared to human experiments, the training and fine-tuning of LLMs demand substantial computational resources, making these processes sometimes cost-prohibitive (Costa 2025). This high computational burden also limits the practical scale of simulations and can hinder the application of classical sensitivity analysis approaches, which are vital for understanding model robustness and parameter influence in traditional social simulations (Piao et al. 2025; Wu et al. 2025). The computational intensity of LLMs means that running extensive simulations, especially those requiring many iterations or large agent populations, can be significantly more resource-intensive than traditional agent-based models. This contrasts sharply with the lower computational demands of many classical ABMs, which can often be run on standard hardware for extensive parameter sweeps and sensitivity analyses.
The combination of the ”black-box” nature of LLMs and their inherent biases presents a significant challenge, particularly when these models are applied in high-stakes social simulations. The opacity of LLM decision making processes (Li 2024), coupled with the fact that their internal processing does not consistently mitigate biases (Kim et al. 2025), creates a situation where the outputs of these simulations could be uninterpretable, inaccurate, or even discriminatory. When these models are used for critical applications like policy testing or urban planning, where direct human validation might be challenging or impossible (Wu et al. 2025), there is a substantial risk of propagating flawed or biased conclusions. This moves beyond a mere technical hurdle; it highlights a critical ethical responsibility for researchers and developers in computational social science. Without robust interpretability tools and comprehensive validation against real-world data, the unexamined nature and embedded biases of LLMs could lead to misinformed policy decisions or exacerbate existing societal inequalities, even when the intentions behind their use are positive.
Beyond these technical and ethical concerns, using an LLM to generate agent behavior fundamentally alters the modeling process itself. Unlike traditional agent-based modeling, where the designer explicitly selects or constructs behavioral models based on theory or empirical evidence, LLM-driven agents bypass this stage: the behavioral model is implicit, embedded within billions of opaque parameters learned from large-scale data. As a result, the modeller relinquishes the ability to clearly define or justify the assumptions underlying agent behavior. This lack of epistemic control is especially problematic in scientific contexts, where the rationale for modeling choices must be transparent and subject to scrutiny. While LLMs may, in theory, reflect statistical patterns closer to real-world data, this advantage is undermined if the underlying mechanisms remain inaccessible and unvalidated.
4.3 LLM Applications: Interactive Use vs. Social Simulation Challenges
In contexts such as educational games, training simulations, or immersive environments (Taillandier et al. 2019b), LLMs offer direct and significant benefits. Here, the primary goal is often to create a ”believable” and ”engaging” experience rather than a perfectly accurate, verifiable replication of reality (Costa 2025). LLMs excel at simulating dynamic, personalized interactions with artificial agents, thereby significantly enhancing user engagement. They can provide adaptive learning pathways, tailoring content and activities to individual needs and preferences. This fosters critical thinking by presenting users with diverse scenarios and responses that encourage problem-solving and decision-making (Hu et al. 2025; Costa 2025). For instance, in a model on teacher-student interactions,LLMs can provide personalized feedback to learners, offering constructive criticism and guidance that is difficult to scale in traditional educational settings. They also contribute to the creation of immersive learning environments through gamification, where learning objectives are interwoven with interactive narratives and challenges (Hu et al. 2025; Costa 2025). The integration of LLMs with virtual and augmented reality holds promise for creating highly immersive and interactive learning experiences, allowing users to step into simulated environments and engage with AI agents as main characters (Costa 2025). Furthermore, LLMs can facilitate personalized learning paths, adapting content and activities to individual needs and preferences, and fostering human-AI collaboration where AI assists teachers with repetitive tasks, freeing them to focus on student-centric activities (Costa 2025).
However, for “pure social simulation” aimed at understanding or predicting complex societal phenomena, the application of LLMs is more complicated due to existing limitations (Chen et al. 2024b). The inherited biases from training data can lead to inaccurate or discriminatory simulation outcomes (Santurkar 2025). The black-box nature makes it difficult to interpret the underlying mechanisms of simulated social phenomena, hindering scientific explanation and theory building (Li 2024). The heavy computational cost limits the use of classical sensitivity analysis approaches, which are crucial for understanding model robustness and parameter influence in traditional social simulations (Wu et al. 2025). The absence of genuine inner psychology, emotions, and intent in LLMs means their simulated behaviors might not accurately reflect the complex decision-making processes of humans in real social contexts (Wang et al. 2025). The ”average persona” problem and lack of behavioral diversity limit the ability to simulate heterogeneous populations and emergent phenomena that arise from individual differences (Wu et al. 2025). Finally, ensuring alignment with real-world patterns, consistency over time, and reproducibility remains a significant challenge, making it difficult to establish the credibility of findings for scientific discovery (Wu et al. 2025). These challenges necessitate careful consideration of the research question and the appropriate level of fidelity required when deploying LLMs for pure social science inquiry.
4.4 Research Fronts and Future Improvements
Future research must prioritize addressing diversity, bias, sycophancy, alienness, and generalization issues in LLM-based simulations (Wu et al. 2025). To enhance realism, developing more diverse and personalized training datasets, including first-person narratives, psychological insights, and cultural experiences, is crucial. Incorporating external motivation structures could enable more human-like decision-making driven by goals, fears, or aspirations, moving beyond mere statistical prediction. Building richer virtual worlds where AI agents compete, collaborate, or survive in dynamic, resource-driven environments could lead to more lifelike decisions and emergent behaviors. New evaluation frameworks are needed to benchmark AI-driven social simulations more systematically. This includes, but is not limited to, metrics assessing psychological alignment with real human behaviors, in order to better calibrate and refine these models for complex societal dynamics. Simulations may be more valuable when focusing on collective patterns rather than individual trajectories, especially given current limitations in individual behavioral diversity (Wu et al. 2025).
Smaller language models (SLMs) are emerging as potentially more suitable and economical for many agentic systems, offering lower latency, reduced memory and computational requirements, and significantly lower operational costs while maintaining adequate task performance in constrained domains (Belcak et al. 2025). This shift could promote more responsible and sustainable AI deployment, addressing some of the environmental and economic concerns associated with large-scale LLMs (Belcak et al. 2025). Integrating LLMs with established theoretical perspectives from psychology, sociology, and decision science, as well as traditional agent-based models (ABMs), offers a powerful toolset for more nuanced and realistic models (Chen et al. 2024b). This hybrid approach allows researchers to combine the generative flexibility of LLMs with the structured analysis capabilities of ABMs. Hybridization can also take the form of architectural compromises — for instance, assigning a single LLM to generate responses for a group of agents, while injecting agent-specific noise or contextual variations to simulate individual behavioral diversity within the simulation. This strategy offers a pragmatic trade-off between computational efficiency and behavioral differentiation, especially in large-scale systems. Continued focus on ethical alignment is critical, especially as LLMs take on complex agentic roles involving decision-making with humans, ensuring that their behaviors align with human values and societal norms (Liu et al. 2024).
5 Conclusion
LLMs have demonstrated impressive capabilities in mimicking human-like conversation and passing complex psychological tests, including the Turing Test and certain Theory of Mind tasks (Jones & Bergen 2025; Kosinski 2023). These capabilities have spurred the development of advanced LLM-driven social simulation frameworks like Generative Agents (Smallville) and AgentSociety (Park 2023; Piao et al. 2025). However, these successes are rooted in sophisticated pattern recognition rather than genuine understanding, leading to significant limitations. These include inherent biases, a black-box nature, high computational costs, a lack of true inner psychology or behavioral diversity, and issues with hallucination and inconsistency (Schramowski 2024; Poeppel & Marantz 2025; Wu et al. 2025; Li 2024). While LLMs show direct utility and promise in interactive simulations and serious games for engagement and personalized learning (Costa 2025; Hu et al. 2025), their application in pure social simulation for accurate prediction or replication of complex human societies remains challenging and requires careful validation and boundary definition (Chen et al. 2024b; Zhou 2024).
The current landscape suggests a fundamental shift in the approach to social simulation. The limitations of pure LLM-based simulations—such as their inherent biases, black-box nature, computational demands, and tendency towards behavioral homogeneity (Wu et al. 2025; Li 2024)—underscore the enduring value of traditional Agent-Based Models (ABMs) for structured environments and rigorous analysis. This has led to a growing interest in integrating LLMs into classical simulation platforms like GAMA (Taillandier et al. 2019a) and NetLogo (Jimenez-Romero et al. 2025; Zhang et al. 2025a; Liu et al. 2025). This approach, often referred to as LLM-augmented ABM, offers a powerful toolset for researchers, allowing for more nuanced, realistic, and comprehensive models of complex systems and human behaviors (Chen et al. 2024b). Such mixed models can strategically leverage the strengths of LLMs, such as their natural language processing capabilities, context-aware processing, and adaptive behaviors, to enrich agent intelligence (Piao et al. 2025). Concurrently, traditional ABM frameworks provide the structured environment, explicit interaction rules, and mechanisms for rigorous analysis and validation (Piao et al. 2025; Wu et al. 2025; Zhang et al. 2025a). Examples of such integration are already emerging, with toolchains integrating LLMs with NetLogo via Python extensions for prompt-driven behavior generation (Jimenez-Romero et al. 2025). A similar integration effort is also underway in the GAMA platform (Nghi et al. 2025), which supports LLM-based behavior scripting through external APIs and Python bridges, allowing modelers to delegate selected decision-making components to language models while preserving the structure and analytical transparency of rule-based modeling. This integration also offers access to a wide set of models, enabling flexible experimentation and comparative evaluation of different LLM architectures within the same modeling environment. This hybrid approach allows modelers familiar with traditional platforms to harness LLM capabilities, fostering the development of models that combine the strengths of both paradigms.
Invoking large language models (LLMs) during agent-based simulations raises significant computational challenges, especially when such calls occur frequently or at scale, potentially limiting their practical usability in large experiments or real-time applications. In response to these constraints, the discussion about smaller language models (SLMs) as the ”future of agentic AI” for efficiency (Belcak et al. 2025) suggests a move towards modularity, where different models (LLM, SLM, rule-based) are chosen for specific sub-tasks based on their strengths and weaknesses. This indicates a paradigm shift in computational social science: instead of LLMs replacing traditional ABMs, they are becoming powerful components within a broader, more sophisticated modeling ecosystem. The future of social simulation will likely involve highly modular, hybrid architectures that strategically combine LLMs for complex, language driven behaviors with traditional rule-based or statistical models for structured interactions, large-scale dynamics, and rigorous analysis, potentially augmented by SLMs for efficiency in specific sub-tasks. This allows researchers to exploit the unique strengths of LLMs while mitigating their inherent limitations, leading to more robust, interpretable, and scientifically valuable social simulations. This also suggests a need for new methodologies to manage the complexity of these hybrid systems and validate their emergent behaviors.

### v2

@@
-
+arXiv:2507.19364v2[cs.AI] 27 Feb 2026
 Abstract
-This position paper examines the use of Large Language Models (LLMs) in social simulation, analyzing both their potential and their limitations from a computational social science perspective.
-The first part reviews recent findings on the ability of LLMs to replicate key aspects of human cognition, including Theory of Mind reasoning and social inference, while also highlighting significant limitations such as cognitive biases, lack of true understanding, and inconsistencies in behavior.
-The second part surveys emerging applications of LLMs in multi-agent simulation frameworks, focusing on system architectures, scale, and validation strategies.
-Notable projects such as Generative Agents (Smallville) and AgentSociety are discussed in terms of their design choices, empirical grounding, and methodological innovations.
+This position paper examines the use of Large Language Models (LLMs) in social simulation, analyzing their potential and limitations from a computational social science perspective.
+We first review recent findings on LLMs’ ability to replicate key aspects of human cognition, including Theory of Mind reasoning and social inference, while identifying persistent limitations such as cognitive biases, lack of grounded understanding, and behavioral inconsistencies.
+We then survey emerging applications of LLMs in multi-agent simulation frameworks, examining system architectures, scalability, and validation strategies.
+Projects such as Generative Agents (Smallville) and AgentSociety are analyzed with respect to their empirical grounding and methodological design.
 Particular attention is given to the challenges of behavioral fidelity, calibration, and reproducibility in large-scale LLM-driven simulations.
-The final section distinguishes between contexts where LLMs, like other “black-box” systems, offer direct value—such as interactive simulations and serious games—and those where their use is more problematic, notably in explanatory or predictive modeling.
-The paper concludes by advocating for hybrid approaches that integrate LLMs into traditional agent-based modeling platforms (GAMA, Netlogo, etc), enabling modelers to combine the expressive flexibility of language-based reasoning with the transparency and analytical rigor of classical rule-based systems.
-keywords: Large Language Models, Social Simulation, Agent-Based Modeling, Human Behavior Modeling
+Finally, we distinguish between contexts where LLM-based agents provide operational value—such as interactive simulations and serious games—and contexts where their use raises epistemic concerns, particularly in explanatory or predictive modeling.
+We argue that hybrid approaches integrating LLMs into established agent-based modeling platforms such as GAMA and NetLogo may offer a promising compromise between expressive flexibility and analytical transparency.
+Building on this analysis, we outline a conceptual research direction termed Hybrid Constitutional Architectures, which proposes a stratified integration of classical agent-based models (ABMs), small language models (SLMs), and LLMs within established platforms such as GAMA and NetLogo.
 1 Introduction
 Large Language Models (LLMs) have emerged as a transformative class of deep learning architectures capable of generating human-like text across a wide range of domains.
-Built upon the transformer architecture, which employs parallelized attention mechanisms rather than sequential processing, these models are pre-trained on massive corpora comprising web data, encyclopedic content, and code repositories (Vaswani 2017; Devlin 2019; Brown et al.
-2020).
-Through unsupervised or self-supervised learning, LLMs acquire syntactic, semantic, and pragmatic language competencies, enabling them to perform a broad array of downstream tasks such as question answering, summarization, translation, and code generation (Tonmoy 2024).
+Built upon the transformer architecture, which employs parallelized attention mechanisms rather than sequential processing, these models are pre-trained on massive corpora comprising web data, encyclopedic content, and code repositories [74, 20, 11].
+Through unsupervised or self-supervised learning, LLMs acquire syntactic, semantic, and pragmatic language competencies, enabling them to perform a broad array of downstream tasks such as question answering, summarization, translation, and code generation [73].
 Importantly, the magnitude and diversity of these training corpora mean that LLMs have effectively absorbed an immense statistical imprint of human language use across cultures, domains, and contexts.
@@
 Subsequently, these models are fine-tuned via instruction tuning, a supervised process in which models learn to follow natural language commands across a wide range of tasks.
-This is typically achieved using curated instruction datasets (Sanh et al.
-2021; Wei et al.
-2021), sometimes augmented with human feedback to better align model outputs with user intent (Ouyang et al.
-2022).
-A critical area of current investigation concerns the capacity of LLMs to simulate human behavior.
-One classical benchmark in this domain is the Turing Test, proposed by Alan Turing in 1950 as the imitation game, which evaluates whether a machine can produce conversational behavior indistinguishable from that of a human interlocutor (Jones & Bergen 2025).
+This is typically achieved using curated instruction datasets [58, 79], sometimes augmented with human feedback to better align model outputs with user intent [53].
+Given this capability to emulate human-like reasoning patterns, LLMs have naturally attracted significant attention within the Agent-Based Modeling (ABM) community.
+However, as their integration into this field accelerates, it is crucial to distinguish between two emerging uses of Large Language Models in computational social science.
+The first involves using LLMs as coding assistants to facilitate the technical implementation of simulation models (e.g., generating NetLogo or GAMA code).
+The second, which is the sole focus of this paper, is the use of LLMs as the cognitive architecture of the agents themselves—often termed ’Generative Agents.’ We examine how LLMs can replace or augment traditional rule-based cognition to simulate human-like decision-making, social interaction, and reasoning, rather than how they assist in the software engineering process.
+Focusing specifically on this role as a cognitive engine, a critical area of current investigation concerns the capacity of LLMs to authentically simulate human behavior.
+One classical benchmark in this domain is the Turing Test, proposed by Alan Turing in 1950 as the imitation game, which evaluates whether a machine can produce conversational behavior indistinguishable from that of a human interlocutor [35].
 Recent empirical findings suggest that state-of-the-art LLMs, notably LLaMa-3.1 and GPT-4.5, have demonstrated convincing performance on a three-party version of this test.
-In these evaluations, LLaMa-3.1 was perceived as human in 56% of cases, while GPT-4.5 achieved a 73% success rate (Jones & Bergen 2025).
+In these evaluations, LLaMa-3.1 was perceived as human in 56% of cases, while GPT-4.5 achieved a 73% success rate [35].
 This represents a notable advancement over previous models, which had only succeeded under less rigorous two-party test conditions.
 These results underscore the impressive ability of modern LLMs to adopt complex linguistic styles and behavioral nuances, particularly when guided by carefully crafted persona prompts — a form of cognitive prompt engineering that remains a highly precise and non-trivial technique requiring expert design and task sensitivity.
-The apparent success of LLMs in the Turing Test has naturally prompted speculation about their suitability for simulating human agents within computational social systems (Park 2023; Anthis et al.
-2025; Kirkeby-Hinrup & Stenseke 2025).
+The apparent success of LLMs in the Turing Test has naturally prompted speculation about their suitability for simulating human agents within computational social systems [54, 4, 39].
 However, the capacity of LLMs to produce believable human-like dialogue is not indicative of genuine understanding or consciousness.
-Rather, their performance reflects advanced statistical pattern recognition — the ability to mimic linguistic structures convincingly — without necessarily modeling the underlying cognitive or emotional states of a human agent (Jones & Bergen 2025).
-The tendency of human evaluators to anthropomorphize fluent language exacerbates this illusion, as observers may attribute intentionality or mental states to outputs that are, in essence, probabilistically generated (Jones & Bergen 2025), something well-known in psychology as the intentional stance (Dennett 1989).
+Rather, their performance reflects advanced statistical pattern recognition — the ability to mimic linguistic structures convincingly — without necessarily modeling the underlying cognitive or emotional states of a human agent [35].
+The tendency of human evaluators to anthropomorphize fluent language exacerbates this illusion, as observers may attribute intentionality or mental states to outputs that are, in essence, probabilistically generated [35], something well-known in psychology as the intentional stance [19].
 This distinction between functional imitation and authentic comprehension is particularly salient for social simulation.
@@
 This position paper proceeds as follows.
-The first section examines the psychological plausibility of LLMs, focusing on their performance in tasks designed to evaluate theory of mind, social reasoning, and behavioral consistency.
-The second section surveys recent efforts to embed LLMs in agent-based social simulations, analyzing architectural strategies, scalability, and validation methodologies.
-The third section discusses both the potential contributions and limitations of LLM-based simulations, with particular emphasis on issues such as bias, interpretability, computational cost, and appropriate domains of application.
-Finally, the last section outlines future directions, including the integration of LLMs into classical modeling environments (e.g., GAMA, NetLogo) to support hybrid models that combine the explanatory clarity of rule-based approaches with the generative flexibility of LLMs.
-2 LLMs in Representing Human Behavior
+Section 2 examines the psychological plausibility of LLMs as individual agents, evaluating their performance in social reasoning tasks while highlighting intrinsic cognitive limitations—such as hallucinations, inconsistency, and statistical biases—that challenge their reliability.
+Section 3 surveys the state of the art in LLM-based social simulation, reviewing emerging platforms, cognitive architectures for multi-agent systems, and current methodological strategies for empirical validation.
+Section 4 critically discusses the methodological barriers to scaling these agents into complex simulations, with particular emphasis on the ”micro-macro validity gap” the loss of epistemic control due to model opacity, and the ethical risks of automation bias.
+Finally, Section 5 proposes a synthesis of future research directions, examining the trade-offs between data-driven behavioral fine-tuning and theory-driven cognitive alignment.
+It advocates for ’Hybrid Constitutional Architectures’— a stratified paradigm that combines Small Language Models (SLMs) with theoretical psychological grounding—and discusses their integration into classical modeling environments (e.g., GAMA, NetLogo) to ensure groundedness and causal validity, particularly in data-scarce and culturally diverse contexts.
+To ensure a comprehensive coverage of this rapidly evolving field, the literature selection for this position paper combines a systematic search with targeted exploratory retrieval.
+First, a structured query was conducted on the Scopus database (cutoff date: February 12) using the following boolean string: TITLE-ABS-KEY ((”Large Language Model*” OR ”Generative Agent*”) AND (”Social Simulation” OR ”Agent-Based”)).
+This initial search yielded a corpus of recent articles (2023-2026), from which we selected seminal works based on citation count and relevance to social science applications.
+Second, given the interdisciplinary and fast-paced nature of the domain (where many breakthroughs appear as preprints), we complemented this systematic extraction with LLM-assisted retrieval (using Gemini 3 Pro).
+This model were used to identify specific niche applications and to uncover cross-disciplinary links (e.g., cognitive psychology references) that keyword-based searches might miss.
+Crucially, human verification remained the final filter: every reference suggested by AI tools was manually reviewed by the authors to ensure scientific rigor, correct attribution, and contextual relevance to the arguments presented.
+2 LLMs in Simulating Human Behavior: Capabilities and Cognitive Limitations
 A critical question in assessing the suitability of LLMs for social simulation lies in evaluating how accurately they capture human psychological, emotional, and behavioral traits.
@@
 This gives rise to a powerful illusion: the model appears to “understand,” when in fact it is statistically interpolating between trillions of human-generated expressions.
-Thus, while LLMs can convincingly mimic language use and pass superficial behavioral tests, the extent to which their outputs align with actual human cognition, reasoning patterns, emotional expression, and social bias remains contested.For instance, recent studies, such as Molchanova et al.
-(2025), raise important concerns about the depth and consistency of LLMs’ psychological realism by exploring their potential to simulate stable and diverse personality traits.
+Thus, while LLMs can convincingly mimic language use and pass superficial behavioral tests, the extent to which their outputs align with actual human cognition, reasoning patterns, emotional expression, and social bias remains contested.For instance, recent studies, such as [50], raise important concerns about the depth and consistency of LLMs’ psychological realism by exploring their potential to simulate stable and diverse personality traits.
 This section provides a structured review of the evidence regarding these capabilities and limitations.
@@
 It is considered a foundational element of social cognition and typically emerges in early childhood.
-One of the canonical tests to assess ToM in developmental psychology is the Sally-Anne test (Baron-Cohen et al.
-1985).
+One of the canonical tests to assess ToM in developmental psychology is the Sally-Anne test [7].
 In this task, a child observes a scene in which Sally places a marble in her basket and leaves the room.
@@
 Recent studies have assessed the capacity of LLMs to pass such false-belief tasks.
-Kosinski (2023) tested multiple LLMs, including GPT-3.5 and GPT-4, on 40 ToM problems designed to mirror psychological benchmarks.
+[40] tested multiple LLMs, including GPT-3.5 and GPT-4, on 40 ToM problems designed to mirror psychological benchmarks.
 While early models performed near chance levels, GPT-4 achieved over 75% accuracy, comparable to that of a 6 or 7-year-old child.
@@
 The robustness of these findings has been explored in follow-up work.
-Jaiswal (2024) tested LLMs on modified ToM tasks involving less prototypical narratives—such as a cat hiding a key or a user renaming folders offline—and found a significant drop in model performance.
+[32] tested LLMs on modified ToM tasks involving less prototypical narratives—such as a cat hiding a key or a user renaming folders offline—and found a significant drop in model performance.
 Performance was highly sensitive to phrasing; LLMs responded more accurately to questions using cognitive verbs (e.g., “think”) than to those phrased with action-oriented verbs (e.g., “look”), suggesting shallow linguistic pattern recognition rather than deep inference.
 In addition to false-belief tasks, LLMs have been evaluated on their ability to interpret irony, indirect speech acts, and social narratives.
-Strachan et al.
-(2024) found that GPT-4 matched or exceeded average human performance in recognizing indirect requests, detecting sarcasm, and attributing communicative intent in narrative contexts.
+[67] found that GPT-4 matched or exceeded average human performance in recognizing indirect requests, detecting sarcasm, and attributing communicative intent in narrative contexts.
 However, these results vary significantly with prompt design and context framing, raising questions about generalizability.
@@
 Emotions play a crucial role in human decision-making and social interaction.
-Unlike rule-based agents or earlier AI systems, LLMs can generate emotionally appropriate language in many contexts (Lee et al.
-2024).
+Unlike rule-based agents or earlier AI systems, LLMs can generate emotionally appropriate language in many contexts [45].
 For example, given a prompt describing a tragic event, GPT-4 can produce empathetic responses and plausible emotional narratives.
@@
 Studies have attempted to evaluate LLMs’ performance on emotional recognition tasks.
-Lai et al.
-(2025) report that LLMs can often classify emotions expressed in text (e.g., anger, sadness) and match tones to situational prompts.
+[42] report that LLMs can often classify emotions expressed in text (e.g., anger, sadness) and match tones to situational prompts.
 Yet these capabilities depend on typical lexical cues and falter when faced with indirect or culturally nuanced expressions.
 Emotional modeling in LLMs remains limited to surface-level mimicry, lacking integration with motivational systems or physiological grounding, both of which are central to human affective behavior.
-2.3 Bias, Representational Gaps, and Cognitive Illusions
-Because LLMs are trained on large-scale human-generated data, they inevitably inherit societal biases embedded in those texts.
-These include gender, racial, and cultural stereotypes.
-Studies have shown that LLMs reproduce associations consistent with biased representations: for instance, associating certain professions with male pronouns or producing racially skewed descriptions of criminal behavior(Santurkar 2025).
-In addition, structural imbalances in data availability mean that voices from underrepresented linguistic or social groups are poorly modeled, resulting in inequitable performance across different populations (Kotek et al.
-2024).
-Beyond representational biases, LLMs also exhibit patterns consistent with cognitive biases observed in humans.(Kim et al.
-2025) found that LLMs replicate anchoring effects, status quo bias, framing effects, and recency bias when posed with structured reasoning tasks.
-Alarmingly, the tendency to apply ”chain-of-thought” prompting—a method designed to enhance logical reasoning—did not consistently reduce these biases.
-In some instances, reasoning chains amplified them, indicating that LLM ”reasoning” may reinforce pre-trained associations rather than correct them.
-Efforts to debias LLMs using prompt engineering or few-shot learning strategies have shown promise but remain partial solutions.
-Schramowski (2024) argues that mitigating bias in LLMs requires not only refined prompting but also upstream control over training data composition and filtering mechanisms.
-2.4 Limitations in Consistency, Hallucination, and Generalization
-A pervasive challenge in LLM behavior is the issue of hallucination, where the model generates outputs that are linguistically fluent but factually incorrect or inconsistent with the input (Tonmoy 2024).
-This phenomenon undermines the credibility of LLM-generated reasoning and introduces substantial risks in simulation contexts.
-Combined with susceptibility to input phrasing and lack of state consistency over time, hallucination further limits the reliability of LLMs for modeling human behavior dynamically.
-Another recurrent issue is inconsistency.
-Identical or nearly identical inputs can elicit divergent responses, even within the same session.
-This instability hinders reproducibility in simulated environments and challenges the interpretability of agent behavior.
-Wu et al.
-(2025) emphasizes that LLM agents, when evaluated across multiple instantiations, display lower variance and reduced behavioral diversity compared to human participants, while also amplifying social stereotypes.
+2.3 Anthropomorphic and Statistical Biases
+2.3.1 Societal and Representational Biases
+Because LLMs are trained on large-scale human-generated data, they inevitably inherit and can perpetuate societal biases embedded in those texts.
+These include explicit and implicit stereotypes related to race, gender, religion, and culture.
+Studies have shown that LLMs reproduce associations consistent with biased representations: for instance, associating certain professions with male pronouns or producing racially skewed descriptions of criminal behavior [60].
+In addition, structural imbalances in data availability mean that voices from underrepresented linguistic or social groups are poorly modeled, resulting in inequitable performance across different populations [41].
+This ”black-box” inheritance of bias persists despite instruction tuning, and differences between model versions can lead to inconsistent user experiences and hidden harms in downstream applications [24].
+2.3.2 Cognitive Biases and the Reasoning Paradox
+Beyond representational biases, LLMs also exhibit patterns consistent with cognitive biases observed in humans—defined as systematic deviations from rationality in judgment.
+[38] found that LLMs replicate anchoring effects, status quo bias, framing effects, and recency bias when posed with structured reasoning tasks.
+While debiasing efforts using prompt engineering have shown promise, they remain partial solutions [62].
+Alarmingly, the tendency to apply ”chain-of-thought” prompting—a method designed to enhance logical reasoning—does not consistently prevent these biases.
+In some instances, reasoning chains actually amplify them, indicating that LLM ”reasoning” may reinforce pre-trained associations rather than correct them [38].
+[62] argues that mitigating these issues requires upstream control over training data composition rather than just refined prompting.
+2.3.3 Convergence towards the ”Average Persona”
+Another bias that raises significant ethical issues is the convergence towards an “average persona.” LLMs fundamentally operate on statistical probabilities, predicting the ”next likely word” based on patterns in their training data rather than deep cognitive understanding [75, 57].
+Consequently, while human decisions are shaped by a rich tapestry of lived experiences and future goals, an LLM merely calculates the most mathematically expected response.
+This results in a flattening of diversity that can inadvertently erase the characteristics of minority subgroups in simulated populations [81, 24].
+Empirical evidence highlights this homogeneity in economic games, where LLMs often produce highly uniform responses compared to the wider range of choices observed in humans [81].
+This trend extends to the political sphere, where models tend to produce narrower distributions of opinions that overrepresent the values of wealthy, young, and politically liberal individuals—the so-called ”WEIRD” demographic bias [59, 29].
+2.3.4 Artifactual Biases from Safety Alignment
+Beyond this statistical convergence, the deliberate ”safety alignment” of LLMs—achieved through Reinforcement Learning from Human Feedback (RLHF)—introduces a further critical risk: the generation of ”false negatives” in exploratory research.
+While training models to be ”helpful and harmless” is essential for commercial products, this normative alignment acts as a systematic filter that ”sanitizes” agent behaviors.
+In the context of social simulation, an over-aligned model may exhibit a ”sycophancy bias” [63], refusing to manifest conflictual, aggressive, or ”dark” social dynamics even when such behaviors are realistic and essential for understanding phenomena like polarization or radicalization.
+Consequently, researchers risk exploring a truncated solution space where catastrophic but plausible social outcomes are suppressed, leading to a dangerous underestimation of systemic risks.
+While context-rich prompts can mitigate this convergence, this variability remains contingent on the user’s intervention rather than an inherent capacity of the model for autonomous diversity [5].
+2.3.5 Bounded Rationality vs. Artifactual Bias
+Finally, a critical theoretical distinction must be made regarding the very definition of ’bias’ in generative social science.
+Following Herbert Simon’s theory of Bounded Rationality [65], human decision-making is inherently constrained by cognitive limitations and heuristics.
+In social simulation, these human deviations from pure logic are not defects to be removed, but constitutive features necessary for realism.
+Therefore, we must rigorously distinguish between ’Desirable Cognitive Biases’—such as loss aversion or emotional reactivity, which reflect authentic human irrationality—and ’Artifactual Model Biases’ stemming from the LLM’s training process.
+The latter include phenomena like ’Sycophancy’ (the tendency to agree with the user to be ’helpful’) [63] or ’Safety Sanitization’ (the refusal to simulate conflictual interactions).
+While Simon’s bounded rationality is a target to be reached, these artifactual biases are ’machine-like’ distortions that flatten social dynamics and must be actively mitigated through architectural constraints.
+2.4 Intrinsic Instability: Hallucinations and Inconsistency
+A pervasive challenge in LLM behavior is the issue of hallucinations, where the model generates outputs that are baseless, untruthful, or diverge from factual reality, even while maintaining a coherent linguistic structure [73].
+These errors often stem from flawed training data or specific architectural biases [48, 31, 28].
+Crucially, when LLMs hallucinate, they typically maintain an ’aura of authority’, generating responses that appear coherent and well-formed.
+This deceptive coherence creates a ’validity illusion’ that can mislead users—especially those lacking domain expertise—and introduces substantial risks in scientific applications where factual accuracy is paramount [64, 68].
+Compounding this issue is the problem of stochastic inconsistency.
+Identical or nearly identical inputs can elicit divergent responses, even within the same session, which hinders reproducibility and challenges the interpretability of agent behavior [23].
+This instability can leave lasting impressions of unreliability, limiting the trustworthiness of the simulation [25].
+Furthermore, while LLMs demonstrate superior performance in unrealistic, ’omniscient’ settings where all agents possess full information [1], they struggle significantly in scenarios characterizing real-world conditions, such as information asymmetry or the management of private knowledge [86].
+Paradoxically, despite this individual instability, Wu et al.
+(2025) emphasize that LLM agents, when evaluated across multiple instantiations, often display lower variance and reduced behavioral diversity compared to human participants, amplifying social stereotypes and tending towards a ’mode collapse’ [81]
 2.5 Summary and Implications
@@
 In the following section, we turn to how these capabilities are being harnessed in practical social simulation frameworks, and how model architecture and evaluation are being addressed in emerging systems.
-3 LLMs in Social Simulation
-The use of LLMs in social simulations builds upon years of research that have seen the application of machine learning techniques in multi-agent systems and, by extension, in agent-based models (Drogoul & Zucker 1998).
+3 State of the Art: LLM-Based Social Simulation
+The use of LLMs in social simulations builds upon years of research that have seen the application of machine learning techniques in multi-agent systems and, by extension, in agent-based models [22].
 It is therefore essential to understand whether LLMs represent a major breakthrough or, more modestly, a logical evolution of this partnership, and to determine if they share the same methodological challenges as previous approaches or introduce new ones.
@@
 We focus on how they are currently used to instantiate artificial societies, the architectural strategies adopted for agent based modeling, and the methodological challenges related to validation and scaling.
-3.1 Foundational and Emerging Platforms
-A foundational illustration of LLM-based social simulation is the Generative Agents framework introduced by Park (2023), widely known as the “Smallville” project.
+3.1 From Generative Agents to Population-Scale Simulation
+A foundational illustration of LLM-based social simulation is the Generative Agents framework introduced by [54], widely known as the “Smallville” project.
 In this system, twenty-five agents inhabit a sandbox environment reminiscent of The Sims, where they autonomously engage in daily routines, form interpersonal relationships, and generate emergent collective behaviors.
@@
 More recent efforts have extended this line of research to address questions of scale, complexity, and empirical grounding.
-One prominent example is AgentSociety, a large-scale platform designed to simulate the dynamics of human societies with over 10,000 LLM-based agents (Piao et al.
-2025).
+One prominent example is AgentSociety, a large-scale platform designed to simulate the dynamics of human societies with over 10,000 LLM-based agents [56].
 Each agent is instantiated using GPT-4 and equipped with cognitive modules for memory, goal management, decision-making, and social relationship tracking.
@@
 Its validation strategy emphasizes empirical fidelity by replicating outcomes observed in behavioral experiments and large-scale social surveys.
-Complementing these large-scale efforts, Simulate Anything offers a flexible platform for generating demographically diverse agent populations using LLMs, leveraging real-world user data for initialization and sampling (Wang 2025).
+Beyond general-purpose social interaction, emerging frameworks are now targeting specific socio-technical systems where physical constraints are paramount.
+Notably, GATSim (Generative Agent Transport Simulation) is a specialized cognitive framework for urban mobility simulation [49].
+Moving away from traditional utility-based discrete choice models, their generative agents demonstrate the ability to plan complex, realistic daily travel itineraries.
+This work highlights a crucial evolution: the application of LLMs is shifting from purely conversational environments (like Smallville) to spatially constrained domains, where agents must balance personal preferences with physical transportation networks .
+Complementing these large-scale efforts, Simulate Anything offers a flexible platform for generating demographically diverse agent populations using LLMs, leveraging real-world user data for initialization and sampling [78].
 The system supports scenario-specific tuning and hypothesis testing across domains such as politics, journalism, and economics.
 Its principal contribution lies in enabling heterogeneous role configurations and behavioral variability, though its validation methodologies are comparatively underdeveloped.
-Another significant contribution in this domain is S3 (Social-network Simulation System), which specifically focuses on social network dynamics through LLM-empowered agents (Gao et al.
-2025).
+Another significant contribution in this domain is S3 (Social-network Simulation System), which specifically focuses on social network dynamics through LLM-empowered agents [26].
 S3 demonstrates particular strengths in simulating three pivotal aspects of human social behavior: emotion, attitude, and interaction behaviors.
@@
 Several more recent platforms have further advanced the field in terms of architectural modularity and methodological rigor.
-GenSim, for instance, is a general-purpose simulation engine that supports tens of thousands of LLM agents (Tang et al.
-2025).
+GenSim, for instance, is a general-purpose simulation engine that supports tens of thousands of LLM agents [72].
 It introduces structured abstractions for social routines, long-term memory management, and adaptive behavior correction, all orchestrated by a dynamic scheduling system.
 Validation combines empirical replication, face validity assessments, and anomaly detection to ensure both plausibility and statistical consistency.
-Another emerging contribution is AgentTorch, a modular and scalable framework that unifies large-scale agent-based modeling with differentiable programming and LLM-powered behavior synthesis (Chopra et al.
-2023).
+Another emerging contribution is AgentTorch, a modular and scalable framework that unifies large-scale agent-based modeling with differentiable programming and LLM-powered behavior synthesis [15].
 Instead of relying on expensive real-time prompting for each agent, AgentTorch uses a small number of archetypal LLM agents to generate compact, reusable policy representations.
@@
 Whereas narrative-first systems emphasize fine-grained agent individuality, AgentTorch focuses on experimental control, statistical robustness, and scientific reproducibility, making it particularly suited for large-scale hypothesis testing and parameter exploration in computational social science.
-SALLMA (Scalable Architecture for LLM Multi-Agent Applications) adopts a layered architecture, separating operational processes (intent formation, task execution, communication) from knowledge-level components (agent profiles, shared memory, workflows) (Becattini et al.
-2025).
+SALLMA (Scalable Architecture for LLM Multi-Agent Applications) adopts a layered architecture, separating operational processes (intent formation, task execution, communication) from knowledge-level components (agent profiles, shared memory, workflows) [8].
 This modular design facilitates agent specialization and integration with external tools via APIs.
 Its validation pipeline combines automated benchmarking with expert-in-the-loop review, aided by interactive dashboards and simulation playback for iterative refinement.
-SocioVerse represents a shift toward population-scale calibration (Zhang et al.
-2025b).
+SocioVerse represents a shift toward population-scale calibration [85].
 It initializes agents from millions of real-world user profiles to mirror authentic distributions of attitudes, preferences, and social structures.
@@
 Its validation strategy emphasizes both empirical benchmarking and participatory evaluation, particularly focusing on the accurate representation of minority groups and on interactive adjustment of simulation parameters.
-Similarly, the LLM-AIDSim framework integrates LLMs (Llama3:8b) into ABM for influence diffusion, enabling language-level responses and deeper insights into user agent interactions (Zhang et al.
-2025a).
+Similarly, the LLM-AIDSim framework integrates LLMs (Llama3:8b) into ABM for influence diffusion, enabling language-level responses and deeper insights into user agent interactions [84].
 Collectively, these platforms reflect the rapid evolution of LLM-based social simulation toward greater scale, demographic realism, and methodological transparency.
@@
 Together, they mark a shift toward simulation systems that not only generate plausible behavior but also aspire to meet the standards of interpretability and scientific robustness necessary for credible computational social science.
-3.2 Architectures for LLM-Based Multi-Agent Systems
-3.2.1 Agent Representation and Cognitive Modules
+3.2 Architecture Patterns for Cognitive Agents
 LLM-based multi-agent models generally instantiate each agent as a distinct LLM instance or session using prompt engineering.
-Architectures typically involve three interconnected modules: a memory system that encodes and retrieves past experiences; a reflection and summarization layer that condenses observations into mental models; and a planning component that uses these reflections to generate actions (Park 2023; Li et al.
-2024).
+Architectures typically involve three interconnected modules: a memory system that encodes and retrieves past experiences; a reflection and summarization layer that condenses observations into mental models; and a planning component that uses these reflections to generate actions [54, 46].
 These modules are often implemented as chained or nested prompts, with each step leveraging the LLM’s generative capacity to simulate cognitive processes such as recall, reasoning, and intention formation.
-For instance, the Humanized Agent-Based Models (h-ABM) framework, proposed by Atchade et al.
-(2024), outlines a modular architecture for integrating LLMs into ABMs to achieve more cognitively and emotionally realistic agents.
-Memory Systems and Psychological Foundations
+For instance, the Humanized Agent-Based Models (h-ABM) framework, proposed by [6], outlines a modular architecture for integrating LLMs into ABMs to achieve more cognitively and emotionally realistic agents.
+3.2.1 Memory Systems and Psychological Foundations
 Many LLM-based agent architectures draw inspiration from cognitive psychology, particularly theories of episodic and semantic memory.
 For example, the memory systems in Generative Agents and AgentSociety are designed to mimic human-like recall, using vector databases or embedding-based retrieval to surface relevant past experiences.
-Some frameworks explicitly reference psychological models, such as the Atkinson-Shiffrin multi-store model or the concept of “theory of mind,” to inform the design of agent cognition and social reasoning (Li et al.
-2024; Doran & et al.
-2023).
+Some frameworks explicitly reference psychological models, such as the Atkinson-Shiffrin multi-store model or the concept of “theory of mind,” to inform the design of agent cognition and social reasoning [46, 21].
 Reflection modules often prompt the LLM to summarize recent events, extract intentions, or infer the mental states of other agents, thus operationalizing constructs from social and developmental psychology.
-Established cognitive architectures such as SOAR (Laird 2012) and ACT-R (Anderson et al.
-2004) have provided theoretical foundations for agent design.
-SOAR, developed by John Laird, Allen Newell, and Paul Rosenbloom at Carnegie Mellon University (Laird 2012), embodies the problem space hypothesis, suggesting that all goal-oriented behavior can be characterized as a search in a problem space defined by states, operators, and goals.
+Established cognitive architectures such as SOAR [43] and ACT-R [3] have provided theoretical foundations for agent design.
+SOAR, developed by John Laird, Allen Newell, and Paul Rosenbloom at Carnegie Mellon University [43], embodies the problem space hypothesis, suggesting that all goal-oriented behavior can be characterized as a search in a problem space defined by states, operators, and goals.
 This cognitive architecture divides knowledge into procedural memory (skills and know-how), semantic memory (facts about the world), and episodic memory (memories of experiences).
-Similarly, ACT-R (Adaptive Control of Thought-Rational) (Anderson et al.
-2004) posits that human cognition can be modeled through interconnected cognitive and perceptual modules that operate according to rational principles.
+Similarly, ACT-R (Adaptive Control of Thought-Rational) [3] posits that human cognition can be modeled through interconnected cognitive and perceptual modules that operate according to rational principles.
 These architectures inform the contemporary design of LLM agents by providing structured frameworks for organizing memory, decision-making, and learning.
-Planning and Decision-Making
+3.2.2 Planning and Decision-Making
 Planning modules in LLM-based multi-agent systems typically exploit the model’s capacity to generate contextually appropriate actions by conditioning outputs on agent-specific goals, environmental affordances, and social context.
@@
 This approach emphasizes the structured design of inputs to steer agents’ reasoning, allowing for complex decision-making without the need for explicit reward modeling.
-A hybrid approach is exemplified by the LGC-MARL framework proposed by Jia et al.
-(2025), which combines an LLM-based planner with graph-structured multi-agent reinforcement learning.
+A hybrid approach is exemplified by the LGC-MARL framework proposed by [33], which combines an LLM-based planner with graph-structured multi-agent reinforcement learning.
 In this architecture, the LLM is tasked with decomposing high-level instructions into executable subtasks, evaluating their coherence and feasibility, and formulating collaborative strategies among agents before final action selection.
-In contrast, more conventional social simulation systems rely predominantly on prompt-engineering strategies to define agent roles, goals, and reasoning pathways, typically without integrating reinforcement signals or dynamic adaptation mechanisms (Junprung 2023).
+In contrast, more conventional social simulation systems rely predominantly on prompt-engineering strategies to define agent roles, goals, and reasoning pathways, typically without integrating reinforcement signals or dynamic adaptation mechanisms [36].
 These approaches underscore the current design trade-off between architectural complexity and ease of interpretability in the planning components of LLM-driven simulations.
-Orchestration and Communication
+3.2.3 Orchestration and Communication
 In LLM-based multi-agent systems, a centralized orchestration layer is commonly adopted to manage simulation time-stepping and regulate agent interaction flows.
@@
 These platforms often support modular ”blocks,” which are reusable behavioral units that allow for systematic experimentation and increased extensibility.
-Integration with external APIs (e.g., web search, databases, calculators) is also crucial, as it enables agents to access additional information and tools, thereby augmenting their reasoning and decision-making capabilities (Chen et al.
-2024b).
+Integration with external APIs (e.g., web search, databases, calculators) is also crucial, as it enables agents to access additional information and tools, thereby augmenting their reasoning and decision-making capabilities [14].
 3.2.4 Specialization and Behavioral Diversity
 LLM-based multi-agent system (MAS) frameworks are progressively integrating sophisticated specialization protocols, allowing individual agents to adopt roles characterized by unique behavioral repertoires.
-This specialization is typically realized through several mechanisms: the application of tailored prompt templates that guide an agent’s responses (e.g., (Wei et al.
-2022)), which can enforce specific communication styles or decision-making heuristics; the implementation of role-specific memory structures that store relevant knowledge and past interactions (e.g., (Park 2023)), enabling agents to leverage historical data pertinent to their designated function; and scenario-driven fine-tuning processes that adapt an agent’s core model to specific operational contexts (e.g., (Ouyang et al.
-2022)), thereby optimizing their performance for particular tasks or environments.
+This specialization is typically realized through several mechanisms: the application of tailored prompt templates that guide an agent’s responses (e.g., [80]), which can enforce specific communication styles or decision-making heuristics; the implementation of role-specific memory structures that store relevant knowledge and past interactions (e.g., [54]), enabling agents to leverage historical data pertinent to their designated function; and scenario-driven fine-tuning processes that adapt an agent’s core model to specific operational contexts (e.g., [53]), thereby optimizing their performance for particular tasks or environments.
 Furthermore, some cutting-edge systems are exploring methods to inject personality traits or demographic attributes into agents.
-This can be achieved either by conditioning prompts with specific personality descriptors (e.g., ”act as a skeptical researcher”) or by sampling attributes from real-world survey distributions (e.g., age, occupation, cultural background), thereby enhancing the realism and diversity of agent behaviors and fostering more nuanced social dynamics (Wang et al.
-2024; Chen et al.
-2024b).
-These innovative techniques clearly enhance existing population synthesis algorithms (e.g., (Chapuis et al.
-2022)) and enable researchers to explore complex emergent behaviors within agent populations, thereby gaining deeper insights into human-like interactions.
-3.3 Validation Strategies and Data Sources
+This can be achieved either by conditioning prompts with specific personality descriptors (e.g., ”act as a skeptical researcher”) or by sampling attributes from real-world survey distributions (e.g., age, occupation, cultural background), thereby enhancing the realism and diversity of agent behaviors and fostering more nuanced social dynamics [77, 14].
+These innovative techniques clearly enhance existing population synthesis algorithms (e.g., [12]) and enable researchers to explore complex emergent behaviors within agent populations, thereby gaining deeper insights into human-like interactions.
+3.3 Current Validation Methodologies
 3.3.1 Empirical Validation and Benchmarking
@@
 Among the most rigorous efforts, frameworks like AgentSociety adopt multi-faceted validation strategies that compare simulated outputs with real-world data.
-Empirical benchmarking typically involves aligning simulation results with social indicators such as voter turnout, economic behavior, or survey-derived attitudinal distributions from established instruments like the General Social Survey or the World Values Survey (Piao et al.
-2025).
+Empirical benchmarking typically involves aligning simulation results with social indicators such as voter turnout, economic behavior, or survey-derived attitudinal distributions from established instruments like the General Social Survey or the World Values Survey [56].
 In addition, some studies replicate classical experimental paradigms—such as public goods games or rumor propagation scenarios—within the simulation environment, and evaluate whether the emergent outcomes converge toward established empirical findings.
 Longitudinal consistency is also a critical focus, particularly in simulations involving extended time horizons.
-Researchers analyze whether agent behavior remains stable and coherent across simulation runs, and whether emergent patterns are reproducible under minor changes to initial conditions or prompt phrasing (Wu et al.
-2025).
+Researchers analyze whether agent behavior remains stable and coherent across simulation runs, and whether emergent patterns are reproducible under minor changes to initial conditions or prompt phrasing [81].
 These approaches aim to assess not only face plausibility but also the robustness and generalizability of the simulated social dynamics.
@@
 Calibration against real-world survey data constitutes another widely used strategy.
-Here, demographic or attitudinal profiles from empirical sources are used to condition LLM agents, and human evaluators subsequently assess the extent to which the simulated agents reflect the intended population characteristics (Chen et al.
-2024b).
-Nevertheless, recent findings suggest that human raters may be influenced by non-epistemic factors—such as assertiveness, linguistic fluency, or the perceived authority of LLM outputs—leading to potential overestimation of their factual or analytical accuracy (Chen et al.
-2024a).
+Here, demographic or attitudinal profiles from empirical sources are used to condition LLM agents, and human evaluators subsequently assess the extent to which the simulated agents reflect the intended population characteristics [14].
+Nevertheless, recent findings suggest that human raters may be influenced by non-epistemic factors—such as assertiveness, linguistic fluency, or the perceived authority of LLM outputs—leading to potential overestimation of their factual or analytical accuracy [13].
 3.3.3 Specialized Validation Methods
 In response to the methodological challenges of evaluating LLM-generated behavior, several specialized validation approaches have emerged.
-Face validation techniques rely on the visualization of agent behavior, often through animation or dashboards, to elicit expert feedback on whether the behavior appears theoretically plausible (Park 2023).
+Face validation techniques rely on the visualization of agent behavior, often through animation or dashboards, to elicit expert feedback on whether the behavior appears theoretically plausible [54].
 Exploratory model behavior analysis is another technique, in which researchers systematically vary input conditions to test whether emergent behaviors align with theoretical expectations.
@@
 Sensitivity analysis further complements these methods by evaluating the robustness of simulation results to perturbations in initial conditions and parameter values.
-One concrete instantiation of this multi-level validation approach is the S3 system (Gao et al.
-2025), which conducts evaluations at both individual and population levels using real-world social network data.
+One concrete instantiation of this multi-level validation approach is the S3 system [26], which conducts evaluations at both individual and population levels using real-world social network data.
 Their validation strategy demonstrates promising accuracy in replicating observed social phenomena, including the propagation of information, attitudes, and emotions across network structures.
-3.3.4 Validation Challenges and Open Issues
-Despite these developments, substantial obstacles remain.
-One pervasive challenge is the convergence of LLMs toward an ”average persona” (Wu et al.
-2025), which can suppress behavioral heterogeneity and thereby reduce the capacity of simulations to model complex social phenomena that rely on diversity of perspective and action.
-In many domains, the lack of clear empirical ground truth complicates efforts to validate simulated phenomena with sufficient rigor.
-Another critical issue concerns temporal instability—agents may exhibit behavioral drift over the course of long simulation runs, compromising reproducibility and longitudinal coherence (Wu et al.
-2025).
-Human evaluation, while informative, introduces its own set of biases.
-Recent findings suggest that human raters may be influenced by non-epistemic factors—such as assertiveness, linguistic fluency, or the perceived authority of LLM outputs—leading to potential overestimation of their factual or analytical accuracy (Steyvers et al.
-2024).
-Moreover, recent studies underscore the limitations of omniscient simulation settings, where agents are given access to complete information.
-While such configurations enhance performance, they are far removed from real-world conditions, which are typically characterized by asymmetrical or incomplete information (Zhou et al.
-2024).
-In light of these challenges, Wu et al.
-(2025) proposes a pragmatic framework for determining the conditions under which LLM-based social simulations can meaningfully advance scientific understanding.
-They argue that these simulations are most defensible when used to explore collective patterns rather than individual decision trajectories, when agent outputs align with population-level averages despite limited within-population variance, and when validation methods are carefully matched to the research objectives and simulation scale.
 3.3.4 Data Sources and Comparative Benchmarks
 The development of standardized benchmarks has been instrumental in advancing the empirical evaluation of LLM agents in social simulations.
-For instance, the SocialIQA dataset (Sap et al.
-2019) evaluates an LLM’s capacity for common sense social reasoning by presenting situations involving physical, social, and emotional causality and testing whether the model can infer likely consequences.
-Another example is CrowS-Pairs (Nangia et al.
-2020), which assesses social bias by contrasting sentence pairs that differ only by protected attributes (e.g., gender or race), allowing systematic measurement of discriminatory tendencies.
-Additionally, benchmarks like BiosBias (De Arteaga et al.
-2019) evaluate demographic fairness by examining whether the model’s predictions or representations vary unjustly depending on demographic factors such as gender or ethnicity.
-While these datasets do not address agent consistency across sessions directly, related work such as Zhang et al.
-(2022) examines the consistency of language models by testing response variability to semantically equivalent paraphrased inputs, providing useful insights into reproducibility concerns in agent behavior.
+For instance, the SocialIQA dataset [61] evaluates an LLM’s capacity for common sense social reasoning by presenting situations involving physical, social, and emotional causality and testing whether the model can infer likely consequences.
+Another example is CrowS-Pairs [51], which assesses social bias by contrasting sentence pairs that differ only by protected attributes (e.g., gender or race), allowing systematic measurement of discriminatory tendencies.
+Additionally, benchmarks like BiosBias [18] evaluate demographic fairness by examining whether the model’s predictions or representations vary unjustly depending on demographic factors such as gender or ethnicity.
+While these datasets do not address agent consistency across sessions directly, related work such as [83] examines the consistency of language models by testing response variability to semantically equivalent paraphrased inputs, providing useful insights into reproducibility concerns in agent behavior.
 Comparative studies between human and LLM behavior further inform the validation landscape.
-The EPITOME battery (Karra et al.
-2024) includes a series of experimental tasks designed to test Theory of Mind capabilities in both humans and LLMs.
+The EPITOME battery [37] includes a series of experimental tasks designed to test Theory of Mind capabilities in both humans and LLMs.
 Results suggest that while LLMs can often match or even exceed human-level performance on certain tasks, they also exhibit systematic limitations, particularly in pragmatic reasoning and inference under uncertainty.
@@
 Future work must focus on developing principled approaches for behavioral diversity injection, cross-run reproducibility, and modular interpretability to ensure these systems can meaningfully contribute to computational social science.
-4 Discussion
-4.1 Strengths of LLMs in Social Simulation
-LLMs offer a compelling alternative to traditional human-participant studies, providing a scalable and cost effective means to explore social scenarios (Anthis et al.
-2025; Piao et al.
-2025).
-They enable the rapid simulation of thousands of interactions, facilitating the study of large-scale social dynamics that would be logistically impractical or prohibitively expensive with human subjects.
-This capability significantly reduces the resource burden associated with recruiting, compensating, and managing human participants, making complex social experiments more accessible (Anthis et al.
-2025; Piao et al.
-2025).
-Furthermore, unlike rigid rule-based simulations, LLMs can sometimes exhibit unexpected emergent behaviors that were not explicitly programmed, potentially offering novel perspectives into complex social phenomena (Anthis et al.
-2025).
-This capacity for spontaneous, human-like interaction can lead to fresh insights into social dynamics that might be difficult to anticipate with more “theoretical” or constrained models.
-While LLMs are fundamentally data-driven—learning behavioral patterns from large-scale corpora—whereas classical cognitive architectures are typically theory-driven, grounded in formal models derived from case studies and psychological theory, both approaches can be leveraged to support exploratory theory building in social simulation, i.e.
-understanding the range of possible effects under specific behavioral assumptions before costly real-world experiments are conducted.
-This makes them valuable for pilot studies where the objective is to uncover possibilities rather than to confirm hypotheses (Anthis et al.
-2025).
-Researchers can then leverage LLMs to identify numerous ways in which social phenomena might manifest, exploring the space of potential outcomes without the need to avoid false positives, which is crucial in early-stage research (Wu et al.
-2025).
-LLMs also enable researchers to conduct ethical investigations into sensitive real-world issues, such as social bias or psychological stress, without subjecting human participants to potentially harmful experimental conditions (Anthis et al.
-2025).
-This provides a safe and controlled environment for exploring complex and ethically challenging scenarios.
-The inherent natural language processing capabilities of LLMs allow generative agents to offer intuitive interfaces for human interaction, making simulations more flexible and their internal processing potentially more explainable (Li et al.
-2024).
-This natural language interface simplifies the design and interpretation of agent behaviors, allowing researchers to define complex roles and interactions using human-readable prompts.
-For instance, LLMs can process and summarize large institutional documents, enabling their integration as proxies for organizational actors within socio-environmental simulations (Zeng et al.
-2025).
-Beyond these general benefits, LLMs offer tangible enhancements to established simulation frameworks.
-Consider, for example, a long-standing epidemiological simulation model like COMOKIT (Gaudou et al.
-2020; Taillandier et al.
-2024), designed to model disease spread and policy interventions.
-Traditionally, agent behaviors in COMOKIT are rule-based, limiting their adaptability to nuanced human responses or contextual variations.
-By integrating LLM-driven agents into such a framework, researchers could enable more realistic behavioral responses.
-For instance, instead of static compliance rates, LLM agents could interpret public health messages, react to evolving social norms, or even engage in simulated disinformation spread, reflecting real-world complexities.
-This allows for the exploration of rich, context-sensitive human dynamics – such as hesitancy towards vaccination campaigns or adherence to social distancing measures – which are challenging to capture with purely rule-based systems.
-LLMs’ ability to process and summarize large institutional documents (e.g., policy papers, news articles) further facilitates the definition of complex agent behaviors, enabling a deeper exploration of socio-environmental interactions and policy implications within existing, well-calibrated models.
-Their ability to emulate context-sensitive reasoning and reflect demographically correlated response patterns highlights their potential for modeling institutional agency and culturally embedded decision-making processes.
-Such models can facilitate new forms of human–AI collaboration in complex domains requiring interpretability, adaptive reasoning, and integration across multiple knowledge sources (Zeng et al.
-2025).
-4.2 Limitations and Points of Vigilance
-The ”black-box” nature of Large Language Models (LLMs), stemming from their intricate architectures with billions of parameters and complex non-linear transformations, creates significant challenges for understanding their internal workings and decision-making processes (Li 2024).
-In general, this lack of transparency poses substantial challenges for interpretability (understanding why an LLM arrives at a particular output becoming a speculative exercise), accountability, and trustworthiness, but it can also contribute to ”automation bias”.
-This bias manifests when modellers, particularly those lacking deep expertise in a given subject matter, excessively rely on LLM-generated information, often without critical evaluation.
-The perceived sophistication and apparent fluency of LLM outputs can effectively lead modelers to over-attribute accuracy and authority to the model, even when its underlying reasoning is obscure or flawed (Cummings 2012).
-So when LLMs are used to drive the behavior of agents in simulations, this ”black-box” issue makes it exceedingly difficult to verify the reliability and validity of the simulation’s results.
-For instance, if an agent’s decision is influenced by an LLM, and that decision is uninterpretable, assessing whether the simulated outcome accurately reflects reality or a desired theoretical model becomes nearly impossible.
-This can undermine the scientific rigor and credibility of the simulation.
-The unverified reliability of LLM-driven simulations also raises serious ethical and regulatory concerns, especially when the outputs of these simulations are intended to inform or directly influence decision-making processes (Li 2024).
-For instance, if a simulation guides policy decisions, resource allocation, or even emergency responses, the inability to fully understand and trust the LLM’s contribution to the simulation’s outcome introduces significant risks.
-Regulators and policymakers require assurances that the tools informing critical decisions are robust, transparent, and accountable, which the ”black-box” nature of LLMs directly contravenes.
-The potential for biased or inexplicable outputs from the LLM to cascade through the simulation, leading to unforeseen and potentially harmful real-world consequences, may necessitate a critical re-evaluation of how these agent-based models are integrated into decision-support systems.
-In terms of ethical issues, a significant consequence of their “black-box” nature is that LLMs inherit and can perpetuate societal biases present in their vast training datasets (Santurkar 2025; Kim et al.
-2025).
-These biases encompass explicit and implicit stereotypes related to race, gender, religion, and health, potentially leading to discriminatory outcomes and inconsistent user experiences (Santurkar 2025; Gallegos et al.
-2024).
-Studies have found pervasive stereotype biases mirroring those in society across various social categories and stereotypes, demonstrating sizable effects on discriminatory decisions (Santurkar 2025).
-Furthermore, LLMs exhibit patterns resembling human cognitive biases (e.g., anchoring, framing, recency bias), which are often not directly visible and can implicitly mislead users (Schramowski 2024; Kim et al.
-2025).
-Cognitive bias, defined as a systematic pattern of deviation from rationality in judgment, arises from information processing mechanisms and can lead to inconsistent decision-making (Schramowski 2024).
-Crucially, recent studies highlight that LLMs’ purported reasoning capabilities do not consistently prevent these cognitive biases; in some cases, applying reasoning even increases vulnerability to certain bias types (Kim et al.
-2025).
-Bias persists in LLMs despite instruction tuning, and differences between model versions can lead to inconsistent user experiences and hidden harms in downstream applications (Gallegos et al.
-2024).
-Another bias that raises significant ethical issues is the convergence towards an “average persona”.
-LLMs fundamentally operate on statistical probabilities, predicting the ”next likely word” based on patterns in their training data (Wang et al.
-2025).
-Their similarities to human language are purely functional, not indicative of deep cognitive understanding or conscious intent (Poeppel & Marantz 2025).
-While humans make decisions influenced by a rich tapestry of lived experiences, financial needs, and future goals, an LLM merely calculates the most expected response based on its training data (Wang et al.
-2025) and LLMs therefore often converge towards an ”average persona”.
-This results in reduced behavioral variance and can inadvertently erase the characteristics of minority subgroups, leading to a lack of diversity in simulated populations (Wu et al.
-2025; Gallegos et al.
-2024).
-Empirical evidence shows LLMs producing highly uniform responses compared to the wider diversity observed in human choices, for instance, in economic games where LLMs consistently choose extreme values while humans exhibit a much wider range (Wu et al.
-2025).
-This homogeneity significantly limits their capacity to accurately simulate complex social dynamics that inherently rely on individual and group heterogeneity (Wu et al.
-2025).
-Furthermore, LLMs tend to produce narrower distributions of political opinions and may overrepresent opinions of wealthy, young, and politically liberal individuals in certain countries (Santurkar et al.
-2023; Gover 2023).
-The way prompts are formulated can either reinforce or mitigate this convergence: generic prompts tend to elicit normative, majority-aligned responses, while more specific or context-rich prompts may allow for greater variation (Argyle et al.
-2023).
-However, this variability remains contingent on the user’s intervention, not on an inherent capacity of the model to generate autonomous diversity.
-As a result, LLMs often fail to replicate the spontaneous variability and unpredictability that characterize real human societies—limiting their effectiveness for simulating pluralistic or emergent behaviors across diverse agent populations.
-LLMs are prone to ”hallucination,” generating content that is baseless, untruthful, or diverges from factual reality, even while maintaining a coherent linguistic structure (Tonmoy 2024).
-Hallucinations can stem from flawed data sources or architectural biases (Li 2023; Huang et al.
-2025; Gautam 2025).
-Crucially, when LLMs hallucinate, they often maintain an aura of authority and credibility, generating responses that appear coherent and well formed, which can mislead users, especially those lacking expertise (Simhi et al.
-2025; Su 2024).
-Such deceptive coherence poses a significant risk in scientific applications where factual accuracy is paramount.
-Additionally, LLMs can exhibit ’inconsistency,’ producing varying or contradictory outputs for very similar (or even identical) inputs, which severely undermines their reliability and trustworthiness for scientific applications (Elazar et al.
-2021).
-This inconsistency can leave lasting impressions of unreliability (Gao et al.
-2022).
-LLMs demonstrate superior performance in unrealistic, ”omniscient” simulation settings where all agents have full information (Aher et al.
-2023).
-However, they struggle significantly in scenarios that more accurately reflect real-world conditions characterized by information asymmetry, where agents possess incomplete or private knowledge (Zhou 2024).
-This fundamental mismatch with real-world information dynamics limits their applicability for accurate social prediction.
-While social simulations, including the ones equipped with LLMs, offer cost-efficiency compared to human experiments, the training and fine-tuning of LLMs demand substantial computational resources, making these processes sometimes cost-prohibitive (Costa 2025).
-This high computational burden also limits the practical scale of simulations and can hinder the application of classical sensitivity analysis approaches, which are vital for understanding model robustness and parameter influence in traditional social simulations (Piao et al.
-2025; Wu et al.
-2025).
-The computational intensity of LLMs means that running extensive simulations, especially those requiring many iterations or large agent populations, can be significantly more resource-intensive than traditional agent-based models.
-This contrasts sharply with the lower computational demands of many classical ABMs, which can often be run on standard hardware for extensive parameter sweeps and sensitivity analyses.
-The combination of the ”black-box” nature of LLMs and their inherent biases presents a significant challenge, particularly when these models are applied in high-stakes social simulations.
-The opacity of LLM decision making processes (Li 2024), coupled with the fact that their internal processing does not consistently mitigate biases (Kim et al.
-2025), creates a situation where the outputs of these simulations could be uninterpretable, inaccurate, or even discriminatory.
-When these models are used for critical applications like policy testing or urban planning, where direct human validation might be challenging or impossible (Wu et al.
-2025), there is a substantial risk of propagating flawed or biased conclusions.
-This moves beyond a mere technical hurdle; it highlights a critical ethical responsibility for researchers and developers in computational social science.
-Without robust interpretability tools and comprehensive validation against real-world data, the unexamined nature and embedded biases of LLMs could lead to misinformed policy decisions or exacerbate existing societal inequalities, even when the intentions behind their use are positive.
-Beyond these technical and ethical concerns, using an LLM to generate agent behavior fundamentally alters the modeling process itself.
-Unlike traditional agent-based modeling, where the designer explicitly selects or constructs behavioral models based on theory or empirical evidence, LLM-driven agents bypass this stage: the behavioral model is implicit, embedded within billions of opaque parameters learned from large-scale data.
+4 Methodological Challenges and Epistemological Risks
+The transition from the technical frameworks described in the previous section to rigorous scientific inquiry reveals a distinct set of hurdles.
+While the platforms reviewed above demonstrate that large-scale LLM simulations are technically feasible, feasibility does not equate to scientific validity.
+The very features that make LLMs attractive—their ability to generate fluent, varied, and human-like text at scale—also introduce profound methodological barriers when moving from exploratory modeling to causal explanation.
+This section critically examines these barriers, focusing on the gap between micro-level believability and macro-level veridicality, the loss of epistemic control due to model opacity, and the systemic risks of bias and cost.
+To guide this analysis, Table 1 provides a synthetic overview of these cognitive, methodological, and technical hurdles, alongside the architectural solutions proposed to overcome them.
+Table 1: Taxonomy of Challenges in LLM-based Social Simulation and Proposed Hybrid Mitigations.
+This table synthesizes the methodological barriers identified in Section 4 and maps them to the architectural solutions proposed in Section 5.
+Category Challenge & Mechanism Impact on Social Simulation Hybrid Mitigation (Sec. 5)
+Cognitive & Behavioral The “Average Persona” Risk
+Convergence toward a statistical mean; suppression of heterogeneity [81].
+Replicant Effect: Loss of population diversity; inability to model polarization or minority behaviors.
+Persona Injection & SLMs: Use specialized Small Language Models (SLMs) fine-tuned on local data rather than generic foundation models.
+Generative Exaggeration
+Tendency to caricature specific roles when prompted away from the mean [52].
+Toxic Caricature: Agents exhibit stereotypical or extreme behaviors rather than nuanced human decision-making.
+Constitutional Filtering: Constrain LLM outputs using a theoretical framework (e.g., TPB) to reject implausible extremes.
+Methodological & Epistemic The “Fluency Fallacy”
+High linguistic coherence masking factual hallucinations [66].
+Surface Realism: Simulations appear valid to human evaluators but lack causal veridicality.
+Groundedness via ABM: Use the ABM environment as the “ground truth” (physics, resources) to validate agent actions.
+Data Scarcity Wall
+Lack of training data for non-WEIRD populations [60].
+Cultural Erasure: Inability to simulate specific groups (e.g., Vietnamese farmers) via pure fine-tuning.
+Structural Alignment: Replace missing data with theoretical rules (Constitutional Architecture) to guide reasoning.
+Technical & Operational Temporal Instability
+Context window limits and stochastic drift over long horizons [81].
+Longitudinal Incoherence: Agents “forget” their history or change personality during long simulations.
+External Memory Modules: Offload long-term memory and state management to the classical ABM structure.
+Computational Bottleneck
+High inference costs preventing large-scale runs.
+No Sensitivity Analysis: Inability to perform rigorous statistical validation (1000+ runs). Stratified Architecture: Delegate routine heuristics to lightweight SLMs; reserve LLMs for complex novelty.
+4.1 The Validity Gap: From Micro-Believability to Macro-Veridicality
+While frameworks for large-scale simulation have emerged, a critical methodological barrier remains regarding empirical validation.
+As argued by [44], validation is currently the central challenge for generative social simulation, as the field lacks standardized protocols to move beyond ’surface realism’.
+Most current validation strategies rely on ’micro-level believability’—assessing whether individual agents respond realistically in isolated contexts [54].
+However, establishing micro-validity does not guarantee ’macro-level veridicality’.
+This gap between individual plausibility and collective accuracy is widened by three systemic failures that only manifest at scale.
+First, the convergence toward an ”average persona” [81] suppresses the behavioral heterogeneity necessary for modeling complex social phenomena.
+In many domains, this results in a mode collapse where diverse agents regress to a uniform mean.
+Second, temporal instability—often exacerbated by context window limitations—compromises longitudinal coherence.
+Agents tend to ’forget’ early instructions or exhibit behavioral drift over long simulation runs, making reproducibility—a cornerstone of scientific inquiry—notoriously difficult [81].
+Third, many simulations rely on omniscient settings where agents access complete information to maintain coherence.
+While this enhances performance, it fundamentally contradicts real-world social dynamics characterized by information asymmetry and private knowledge [87].
+Relying on human evaluation to bridge this gap is perilous.
+Recent findings suggest that human raters are influenced by non-epistemic factors—such as the assertiveness and linguistic fluency of LLM outputs—leading to a potential overestimation of their factual accuracy [66].
+This creates a ’validity illusion’: the simulation looks realistic because the text is fluent, but the stochastic and ”black-box” nature of LLMs makes calibration—the standard process of tuning ABM parameters to match observed outputs—notoriously difficult.
+Consequently, until quantitative calibration becomes a primary focus, large-scale LLM simulations remain in the realm of ’plausible storytelling’ rather than predictive science [44].
+In light of these challenges, [81] proposes a pragmatic framework for determining the conditions under which LLM-based social simulations can meaningfully advance scientific understanding.
+They argue that these simulations are most defensible when used to explore collective patterns rather than individual decision trajectories, provided that agent outputs align with population-level averages and that validation methods are carefully matched to the research objectives and simulation scale.
+4.2 The ”Black-Box” Problem and Interpretability
+The ”black-box” nature of Large Language Models, stemming from architectures with billions of parameters and complex non-linear transformations, creates significant challenges for understanding their decision-making processes [47].
+This opacity does not merely hinder interpretability; it actively contributes to ”automation bias.” This phenomenon manifests when modelers, seduced by the sophistication and apparent fluency of LLM outputs, over-attribute accuracy and authority to the model, even when its underlying reasoning is obscure or flawed [17].
+This combination of technical opacity and misplaced human trust becomes dangerous in high-stakes social simulations.
+Because the model’s internal processing does not consistently mitigate biases [38], outputs can be uninterpretable, inaccurate, or discriminatory.
+When such simulations are used for critical applications—like policy testing, resource allocation, or urban planning—where direct human validation is often impossible [81], there is a substantial risk of propagating flawed conclusions.
+As noted by [47], regulators require assurances that decision-support tools are robust and accountable—requirements that the ”black-box” nature of LLMs directly contravenes.
+Without robust interpretability tools, the unexamined nature of these models could lead to misinformed policy decisions or exacerbate societal inequalities, even when the researchers’ intentions are positive.
+Beyond these ethical concerns, using an LLM to generate agent behavior fundamentally alters the scientific modeling process itself.
+Unlike traditional Agent-Based Modeling (ABM), where the designer explicitly constructs behavioral rules based on theory or empirical evidence, LLM-driven agents bypass this stage.
+The behavioral model becomes implicit, embedded within billions of opaque parameters learned from massive datasets.
 As a result, the modeller relinquishes the ability to clearly define or justify the assumptions underlying agent behavior.
@@
 While LLMs may, in theory, reflect statistical patterns closer to real-world data, this advantage is undermined if the underlying mechanisms remain inaccessible and unvalidated.
-4.3 LLM Applications: Interactive Use vs. Social Simulation Challenges
-In contexts such as educational games, training simulations, or immersive environments (Taillandier et al.
-2019b), LLMs offer direct and significant benefits.
-Here, the primary goal is often to create a ”believable” and ”engaging” experience rather than a perfectly accurate, verifiable replication of reality (Costa 2025).
-LLMs excel at simulating dynamic, personalized interactions with artificial agents, thereby significantly enhancing user engagement.
-They can provide adaptive learning pathways, tailoring content and activities to individual needs and preferences.
-This fosters critical thinking by presenting users with diverse scenarios and responses that encourage problem-solving and decision-making (Hu et al.
-2025; Costa 2025).
-For instance, in a model on teacher-student interactions,LLMs can provide personalized feedback to learners, offering constructive criticism and guidance that is difficult to scale in traditional educational settings.
-They also contribute to the creation of immersive learning environments through gamification, where learning objectives are interwoven with interactive narratives and challenges (Hu et al.
-2025; Costa 2025).
-The integration of LLMs with virtual and augmented reality holds promise for creating highly immersive and interactive learning experiences, allowing users to step into simulated environments and engage with AI agents as main characters (Costa 2025).
-Furthermore, LLMs can facilitate personalized learning paths, adapting content and activities to individual needs and preferences, and fostering human-AI collaboration where AI assists teachers with repetitive tasks, freeing them to focus on student-centric activities (Costa 2025).
-However, for “pure social simulation” aimed at understanding or predicting complex societal phenomena, the application of LLMs is more complicated due to existing limitations (Chen et al.
-2024b).
-The inherited biases from training data can lead to inaccurate or discriminatory simulation outcomes (Santurkar 2025).
-The black-box nature makes it difficult to interpret the underlying mechanisms of simulated social phenomena, hindering scientific explanation and theory building (Li 2024).
-The heavy computational cost limits the use of classical sensitivity analysis approaches, which are crucial for understanding model robustness and parameter influence in traditional social simulations (Wu et al.
-2025).
-The absence of genuine inner psychology, emotions, and intent in LLMs means their simulated behaviors might not accurately reflect the complex decision-making processes of humans in real social contexts (Wang et al.
-2025).
-The ”average persona” problem and lack of behavioral diversity limit the ability to simulate heterogeneous populations and emergent phenomena that arise from individual differences (Wu et al.
-2025).
-Finally, ensuring alignment with real-world patterns, consistency over time, and reproducibility remains a significant challenge, making it difficult to establish the credibility of findings for scientific discovery (Wu et al.
-2025).
-These challenges necessitate careful consideration of the research question and the appropriate level of fidelity required when deploying LLMs for pure social science inquiry.
-4.4 Research Fronts and Future Improvements
-Future research must prioritize addressing diversity, bias, sycophancy, alienness, and generalization issues in LLM-based simulations (Wu et al.
-2025).
-To enhance realism, developing more diverse and personalized training datasets, including first-person narratives, psychological insights, and cultural experiences, is crucial.
-Incorporating external motivation structures could enable more human-like decision-making driven by goals, fears, or aspirations, moving beyond mere statistical prediction.
-Building richer virtual worlds where AI agents compete, collaborate, or survive in dynamic, resource-driven environments could lead to more lifelike decisions and emergent behaviors.
-New evaluation frameworks are needed to benchmark AI-driven social simulations more systematically.
-This includes, but is not limited to, metrics assessing psychological alignment with real human behaviors, in order to better calibrate and refine these models for complex societal dynamics.
-Simulations may be more valuable when focusing on collective patterns rather than individual trajectories, especially given current limitations in individual behavioral diversity (Wu et al.
-2025).
-Smaller language models (SLMs) are emerging as potentially more suitable and economical for many agentic systems, offering lower latency, reduced memory and computational requirements, and significantly lower operational costs while maintaining adequate task performance in constrained domains (Belcak et al.
-2025).
-This shift could promote more responsible and sustainable AI deployment, addressing some of the environmental and economic concerns associated with large-scale LLMs (Belcak et al.
-2025).
-Integrating LLMs with established theoretical perspectives from psychology, sociology, and decision science, as well as traditional agent-based models (ABMs), offers a powerful toolset for more nuanced and realistic models (Chen et al.
-2024b).
-This hybrid approach allows researchers to combine the generative flexibility of LLMs with the structured analysis capabilities of ABMs.
-Hybridization can also take the form of architectural compromises — for instance, assigning a single LLM to generate responses for a group of agents, while injecting agent-specific noise or contextual variations to simulate individual behavioral diversity within the simulation.
-This strategy offers a pragmatic trade-off between computational efficiency and behavioral differentiation, especially in large-scale systems.
-Continued focus on ethical alignment is critical, especially as LLMs take on complex agentic roles involving decision-making with humans, ensuring that their behaviors align with human values and societal norms (Liu et al.
-2024).
+4.3 Systemic Bias and the ”Average Persona” Risk
+While Section 2.3 highlighted the intrinsic biases of LLMs, their impact scales non-linearly when applied to multi-agent populations.
+The most critical risk for social simulation is systemic homogenization.
+As noted by [81], the statistical nature of LLMs drives them to converge toward a generic, safe, and ”WEIRD” behavioral mean.
+In a simulation context, this creates a ”Replicant Effect”: even when agents are assigned distinct demographic prompts, their actual decision-making logic collapses into a uniform mode, effectively erasing minority behaviors and extreme viewpoints.
+Moreover, this homogenization is paradoxically accompanied by ”Generative Exaggeration”.
+As demonstrated by [52], when LLMs are prompted to adopt specific personas to escape the ”average,” they often drift into caricature, amplifying stereotypical traits to the point of toxicity rather than maintaining behavioral nuance.
+This suggests that current models struggle to find a middle ground between the ”flatness” of the average persona and the ”parody” of specific roles.
+This ”Average Persona” problem is exacerbated by the ”Data Scarcity Wall.” As discussed regarding the limitations of fine-tuning, simulating specific, under-represented populations—such as smallholder farmers in the Global South—is often impossible because the massive, digitized behavioral logs required for training simply do not exist.
+Consequently, relying on generic LLMs to simulate these groups risks replacing missing data with hallucinated stereotypes.
+To mitigate this, future research must move beyond mere prompt engineering.
+Developing more diverse and personalized training datasets is crucial, specifically incorporating first-person narratives, oral histories, and qualitative psychological insights to capture the ”long tail” of human behavior.
+Furthermore, incorporating external motivation structures (e.g., BDI architectures) can help force agents to deviate from the statistical mean by driving decisions through explicit goals and fears rather than mere text prediction.
+Without these structural interventions, simulations will remain valuable for exploring collective patterns of the ”average” human [81], but will fail to capture the diversity essential for understanding complex societal dynamics.
+4.4 Computational Constraints and Scalability
+While LLMs offer cost-efficiency compared to human experiments, they introduce a drastic computational bottleneck compared to traditional Agent-Based Models (ABMs).
+The resource intensity of LLMs—both in fine-tuning and real-time inference—limits practical scalability [16].
+Crucially, this high cost is not merely a logistical hurdle but a methodological barrier: it hinders the application of classical sensitivity analyses, which require running simulations thousands of times to assess parameter influence and model robustness [56, 81].
+Unlike lightweight rule-based agents that permit exhaustive parameter sweeps on standard hardware, the computational latency of LLMs currently forces researchers to trade statistical rigor for behavioral complexity, often preventing the extensive replications necessary to validate stochastic models.
+4.5 Ethics in Application: Serious Games vs. Predictive Science
+Given the methodological barriers outlined above, it is crucial to delineate the appropriate domains of application for LLM-based agents.
+4.5.1 Interactive Environments and Serious Games
+In contexts such as educational games and training simulations [71], LLMs offer direct and significant benefits.
+Here, the primary goal is often to create a ”believable” and ”engaging” experience rather than a perfectly verifiable replication of social reality [16].
+LLMs excel at simulating dynamic, personalized interactions, enhancing user engagement through adaptive learning pathways and gamified narratives [30].
+However, as highlighted by recent critiques, this application requires a shift from cautious optimism to critical vigilance.
+The primary risk lies in ’negative training transfer’: unlike traditional rule-based simulators where every outcome is verified, an LLM-based agent can hallucinate factual errors with high confidence.
+Recent studies highlight that this ’authoritative misinformation’ can lead to superficial learning and cognitive over-reliance [88].
+Furthermore, instructional quality is threatened by inherent biases; as demonstrated by [76], models often perpetuate harmful stereotypes in personalized content.
+Consequently, the deployment of LLMs in training necessitates rigorous ’red-teaming’ and grounded architectures (RAG) to ensure the simulation remains a tool for learning, not confusion.
+4.5.2 From Prediction to Exploratory Modeling
+For ”pure social simulation” aimed at scientific inquiry, the implications of the ”validity gap” and ”epistemic opacity” (discussed in Sections 4.1 and 4.2) are decisive.
+Because of the risks of hallucination and the ”average persona” bias, we argue that LLMs are currently ill-suited for confirmatory empirical research or precise forecasting (e.g., predicting specific election outcomes or policy impacts).
+Instead, their most scientifically robust application lies in Exploratory Modeling [14].
+In this paradigm, LLMs serve not as oracles of truth, but as engines for hypothesis generation.
+They allow researchers to rapidly explore a vast space of possible social interactions and identify emergent patterns that—while not necessarily predictive of the real world—reveal the logical consequences of specific behavioral assumptions.
+As [81] suggests, these simulations should be viewed as ”reasoning pumps” to test collective dynamics under diverse narrative constraints, rather than as digital twins of reality.
+5 Towards Hybrid Constitutional Architectures
+The current landscape suggests a fundamental shift in the approach to social simulation.
+While LLMs offer compelling opportunities—enabling the rapid simulation of thousands of interactions and uncovering emergent behaviors [4, 56]—the limitations of ”off-the-shelf” models underscore the enduring value of traditional Agent-Based Models (ABMs).
+5.1 Integrating LLMs into Classical Frameworks
+The future of social simulation does not lie in ”off-the-world” AI agents, but in the seamless integration of LLMs within structured platforms such as GAMA [70] and NetLogo [34].
+This approach, which we define as LLM-augmented ABM, addresses the fundamental problem of groundedness: while LLMs manipulate abstract concepts, the ABM provides the ”world model”—the physical, spatial, and resource-based constraints—that dictates the actual viability of AI-proposed actions.
+This synergy allows researchers to move beyond purely narrative modeling.
+Recent work, such as Gao et al.
+(2025) on the S3 system, demonstrates how individual behavior alignment can be synchronized with complex social network structures, ensuring that LLM expressiveness does not sacrifice macroscopic coherence [26].
+Simultaneously, frameworks like AgentTorch [15] open the door to massive scalability through GPU acceleration, enabling simulations of entire populations where the ABM manages the physics of interaction while the LLM simulates context-sensitive psychology.
+For instance, in an epidemiological model like COMOKIT [27, 69], an LLM could interpret evolving social norms or disinformation, while the ABM maintains the structural rigor of disease spread.
+This hybrid necessity is even more acute in high-stakes environments.
+[82] recently demonstrated this in disaster evacuation simulations: while the ABM manages the physical constraints of crowd density and velocity to prevent unrealistic movements, the LLM-enhanced agents simulate the cognitive panic and decision-making under uncertainty.
+This ”Physics-Cognition” split ensures that the simulation remains physically valid while capturing the psychological volatility of human crowds.
+5.2 Cognitive Alignment: Fine-Tuning and its Limits
+To correct the instability of generic models, a promising path lies in behavioral alignment.
+Binz et al.
+(2025) have demonstrated that it is possible to transform an LLM into a cognitively aligned agent (the Centaur model) by training it on massive datasets derived from human psychological experiments.
+This process allows the AI to replicate human learning curves and decision-making biases correlated with human neural signals [10].
+However, this data-driven path hits a critical ”Data Scarcity Wall” as soon as one moves beyond Western (WEIRD) contexts.
+As highlighted by [60], when representative training data is absent, LLMs tend to fall back on generalized stereotypes to fill the gaps.
+Consequently, in social simulation where identity is relational and cultural, modeling the decisions of smallholder farmers in Vietnam is impossible through simple fine-tuning: the necessary digitized behavioral logs do not exist, leaving the model to rely on hallucinated or westernized proxies.
+Pursuing a strategy exclusively based on data risks reinforcing the ”Average Avatar Bias”: an AI that perfectly mimics a mean Western human but fails entirely to represent the specificities of marginalized or culturally distinct groups.
+To break this impasse, alignment must shift from being purely statistical (data-driven) to being structural (theory-driven), justifying the transition toward ”constitutional” architectures.
+5.3 Proposed Paradigm: Hybrid Constitutional Architectures
+We argue that cognitive alignment in social simulation cannot remain purely data-driven; it must be structurally anchored in psychological theory.
+We propose the adoption of Hybrid Constitutional Architectures, a paradigm shift where the LLM’s generative freedom is bounded by a theoretical ’scaffold.’
+In this stratified cognitive process, we move beyond the monolithic agent:
+  • Heuristic Stratum (SLMs): High-frequency, routine behaviors (e.g., daily navigation, basic consumption) are delegated to Small Language Models (SLMs).
+This ensures computational scalability [9] and allows for ’local alignment’—where smaller models are fine-tuned on specific ethnographic data that would be diluted in larger architectures.
+  • Reasoning Stratum (LLM as Proposal Engine): Larger models are invoked only for complex, non-routine social reasoning.
+Crucially, the LLM does not execute the final decision; it acts as a ’Proposal Engine’, generating multiple plausible behavioral trajectories based on the current social context.
+  • The Theoretical Constitution: Every proposal is passed through a validation filter grounded in established frameworks, such as the Theory of Planned Behavior (TPB) [2] or Bounded Rationality principles [65].
+For instance, a proposal that contradicts an agent’s established ’Subjective Norms’ or ’Perceived Behavioral Control’ is rejected or penalized.
+The constitutional layer must itself be treated as a model component subject to audit, since poorly specified normative constraints may reduce harmful outputs while simultaneously suppressing legitimate social heterogeneity.
+  • External Memory Modules (ABM State): To resolve the Temporal Instability inherent to finite context windows, the architecture offloads long-term state (biography, resources, social graph) to the classical ABM environment.
+Rather than relying on the LLM to ’remember’ the entire history, relevant data is retrieved dynamically via Retrieval-Augmented Generation (RAG).
+This separation ensures that agents maintain consistent identities over thousands of simulation steps [55, 54].
+To translate this conceptual architecture into an explicit implementation protocol, we formalize the decision-making process as a stratified execution loop that strictly arbitrates between the ABM, SLMs, and LLMs.
+Let $E_{t}$ denote the global environment state at simulation step $t$, managed by the classical ABM (e.g., GAMA, NetLogo) acting as the absolute ground truth, and $S_{i,t}$ the internal state of agent $i$.
+Arbitration is governed by a context-shift function, $\Delta(E_{t},S_{i,t})$, which continuously evaluates environmental and social stability.
+If $\Delta(E_{t},S_{i,t})\leq\tau$ (where $\tau$ is a predefined stability threshold), the environment is considered stable.
+The agent defaults to the computationally lightweight Heuristic Stratum, generating a routine action via a Small Language Model or classical heuristics: $a_{\text{heur}}=\text{SLM}(E_{t},S_{i,t})$.
+However, if the ABM detects a novel event, a social anomaly, or a critical threshold breach (e.g., a sudden policy change or resource depletion), $\Delta(E_{t},S_{i,t})>\tau$.
+Decision-making is then escalated to the Reasoning Stratum.
+Here, the LLM acts as a proposal engine, generating an intended behavioral trajectory based on context and retrieved memory ($M_{i}$): $a_{\text{prop}}=\text{LLM}(E_{t},S_{i,t},M_{i})$.
+Crucially, this generated intention is never executed directly.
+It must pass through the Theoretical Constitution, formalized as a validation filter $\Phi(a,E_{t},S_{i,t})\in\{0,1\}$.
+Grounded in psychological frameworks like the Theory of Planned Behavior (TPB), $\Phi$ evaluates both the physical possibility of the action within the ABM and its normative plausibility for the specific agent.
+The final action execution follows a strict fallback mechanism:
+$a_{\text{final}}=\begin{cases}a_{\text{prop}}&\text{if }\Delta>\tau\text{ and }\Phi(a_{\text{prop}},E_{t},S_{i,t})=1\\ a_{\text{heur}}&\text{otherwise (fallback or routine)}\end{cases}$ (1)
+Finally, the ABM gathers all validated actions ($\mathbf{A}$), resolves physical conflicts, and updates the environment state for the next step: $E_{t+1}=f_{\text{ABM}}(E_{t},\mathbf{A})$.
+This formal protocol (summarized in Algorithm 1) ensures that the generative power of LLMs is rigorously bounded by causal and computational constraints.
+Algorithm 1 Hybrid Constitutional Execution Loop (Single Agent Step)
+1:Current environment state $E_{t}$, Agent internal state $S_{i,t}$, Stability threshold $\tau$
+2:Updated environment state $E_{t+1}$
+[…]
+[…]
+5:if $\delta>\tau$ then $\triangleright$ Escalate to LLM Reasoning
+6:  $M_{i}\leftarrow\text{RAG}(S_{i,t})$ $\triangleright$ Retrieve contextual memory
+7:  $a_{\text{prop}}\leftarrow\text{LLM}(E_{t},S_{i,t},M_{i})$ $\triangleright$ Generate behavioral proposal
+8:  if $\Phi(a_{\text{prop}},E_{t},S_{i,t})==1$ then $\triangleright$ Constitutional Validation
+9:   $a_{\text{final}}\leftarrow a_{\text{prop}}$
+10:  else
+11:   $a_{\text{final}}\leftarrow a_{\text{heur}}$ $\triangleright$ Fallback due to constitutional rejection
+12:  end if
+13:else
+14:  $a_{\text{final}}\leftarrow a_{\text{heur}}$ $\triangleright$ Default to routine behavior
+15:end if
+16:Execute $a_{\text{final}}$ via ABM engine
+[…]
+To measure the analytical improvement of this hybrid approach against both classical rule-based ABMs and pure LLM-driven simulations, we propose a tripartite evaluation pipeline.
+First, Counterfactual Robustness: unlike classical ABMs that fail outside pre-coded rules, and pure LLMs that may hallucinate physically impossible actions, the hybrid model must be tested against unseen socio-environmental shocks to ensure responses are both socially plausible and physically grounded.
+Second, Behavioral Diversity: since pure LLMs tend to collapse into an ’Average Persona,’ the hybrid architecture—particularly through locally fine-tuned SLMs—must be quantitatively evaluated on its ability to preserve statistical heterogeneity and minority behaviors.
+Finally, Computational Scalability: a key metric of success is the ability to run exhaustive sensitivity analyses (e.g., Monte Carlo simulations with 1000+ runs).
+While pure LLM models face prohibitive API latency and costs, the stratified delegation to SLMs must demonstrate a significant reduction in computational overhead, restoring the rigorous statistical validation expected in traditional social simulation.
+By treating the LLM as a creative engine constrained by a formal psychological constitution, we solve the ’black-box’ problem of intent.
+Decision transparency is regained: the modeler can verify if a choice was made due to a social norm (rule-based) or a creative adaptation (LLM-based).
+This synergy moves the field from ’surface realism’ to causal plausibility, enabling simulations that remain robust even when empirical data is scarce.
+Rather than replacing classical agent-based modeling, Hybrid Constitutional Architectures reposition LLMs as bounded cognitive modules within a transparent causal framework, thereby reconciling expressive flexibility with epistemic accountability.
 6 Conclusion
-LLMs have demonstrated impressive capabilities in mimicking human-like conversation and passing complex psychological tests, including the Turing Test and certain Theory of Mind tasks (Jones & Bergen 2025; Kosinski 2023).
-These capabilities have spurred the development of advanced LLM-driven social simulation frameworks like Generative Agents (Smallville) and AgentSociety (Park 2023; Piao et al.
-2025).
-However, these successes are rooted in sophisticated pattern recognition rather than genuine understanding, leading to significant limitations.
-These include inherent biases, a black-box nature, high computational costs, a lack of true inner psychology or behavioral diversity, and issues with hallucination and inconsistency (Schramowski 2024; Poeppel & Marantz 2025; Wu et al.
-2025; Li 2024).
-While LLMs show direct utility and promise in interactive simulations and serious games for engagement and personalized learning (Costa 2025; Hu et al.
-2025), their application in pure social simulation for accurate prediction or replication of complex human societies remains challenging and requires careful validation and boundary definition (Chen et al.
-2024b; Zhou 2024).
-The current landscape suggests a fundamental shift in the approach to social simulation.
-The limitations of pure LLM-based simulations—such as their inherent biases, black-box nature, computational demands, and tendency towards behavioral homogeneity (Wu et al.
-2025; Li 2024)—underscore the enduring value of traditional Agent-Based Models (ABMs) for structured environments and rigorous analysis.
-This has led to a growing interest in integrating LLMs into classical simulation platforms like GAMA (Taillandier et al.
-2019a) and NetLogo (Jimenez-Romero et al.
-2025; Zhang et al.
-2025a; Liu et al.
-2025).
-This approach, often referred to as LLM-augmented ABM, offers a powerful toolset for researchers, allowing for more nuanced, realistic, and comprehensive models of complex systems and human behaviors (Chen et al.
-2024b).
-Such mixed models can strategically leverage the strengths of LLMs, such as their natural language processing capabilities, context-aware processing, and adaptive behaviors, to enrich agent intelligence (Piao et al.
-2025).
-Concurrently, traditional ABM frameworks provide the structured environment, explicit interaction rules, and mechanisms for rigorous analysis and validation (Piao et al.
-2025; Wu et al.
-2025; Zhang et al.
-2025a).
-Examples of such integration are already emerging, with toolchains integrating LLMs with NetLogo via Python extensions for prompt-driven behavior generation (Jimenez-Romero et al.
-2025).
-A similar integration effort is also underway in the GAMA platform (Nghi et al.
-2025), which supports LLM-based behavior scripting through external APIs and Python bridges, allowing modelers to delegate selected decision-making components to language models while preserving the structure and analytical transparency of rule-based modeling.
-This integration also offers access to a wide set of models, enabling flexible experimentation and comparative evaluation of different LLM architectures within the same modeling environment.
-This hybrid approach allows modelers familiar with traditional platforms to harness LLM capabilities, fostering the development of models that combine the strengths of both paradigms.
-Invoking large language models (LLMs) during agent-based simulations raises significant computational challenges, especially when such calls occur frequently or at scale, potentially limiting their practical usability in large experiments or real-time applications.
-In response to these constraints, the discussion about smaller language models (SLMs) as the ”future of agentic AI” for efficiency (Belcak et al.
-2025) suggests a move towards modularity, where different models (LLM, SLM, rule-based) are chosen for specific sub-tasks based on their strengths and weaknesses.
-This indicates a paradigm shift in computational social science: instead of LLMs replacing traditional ABMs, they are becoming powerful components within a broader, more sophisticated modeling ecosystem.
-The future of social simulation will likely involve highly modular, hybrid architectures that strategically combine LLMs for complex, language driven behaviors with traditional rule-based or statistical models for structured interactions, large-scale dynamics, and rigorous analysis, potentially augmented by SLMs for efficiency in specific sub-tasks.
-This allows researchers to exploit the unique strengths of LLMs while mitigating their inherent limitations, leading to more robust, interpretable, and scientifically valuable social simulations.
-This also suggests a need for new methodologies to manage the complexity of these hybrid systems and validate their emergent behaviors.
+LLMs have demonstrated impressive capabilities in mimicking human-like conversation and passing complex psychological tests [35, 40].
+Yet, as this paper has explored, these successes are often rooted in sophisticated pattern recognition rather than genuine understanding.
+This leads to the ”Fluency Fallacy”: a state where models maintain high linguistic coherence even when hallucinating social facts, creating a ”surface realism” that can mislead researchers into mistaking narrative plausibility for causal validity.
+The transition from individual generative agents to robust social simulations requires moving beyond the Turing Test toward stricter criteria: Behavioral Diversity (to avoid mode collapse towards an average persona) and Counterfactual Robustness (the ability to react realistically to unseen shocks) [81, 44].
+While LLMs show direct utility in interactive simulations and serious games for engagement [16], their application in pure social science demands a critical re-evaluation of epistemic control.
+Ultimately, the future of social simulation does not lie in a choice between the data-driven flexibility of LLMs and the theory-driven rigor of ABM, but in their synergy.
+By coupling the narrative transparency of Generative AI with the structural rigor of Hybrid Constitutional Architectures, we can achieve simulations that are not only plausible but causally valid and intuitively explainable.
+This hybrid paradigm ensures that even in data-scarce environments, artificial societies can serve as meaningful laboratories for understanding the deep-seated nuances of human decision-making.
+Conflict of Interest Statement
+The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
+Author Contributions
+P.T.
+conceived the original structure and wrote the first draft of the manuscript.
+J.D.Z., A.G., B.G., N.Q.H., and A.D.
+substantially contributed to the conceptualization, provided critical feedback, and improved the manuscript.
+The development of this article was a truly collective effort.
+All authors discussed the concepts, read, and approved the final manuscript.